### v1

# From Large to Super-Tiny: End-to-End Optimization for Cost-Efficient LLMs

Abstract
In recent years, Large Language Models (LLMs) have significantly advanced artificial intelligence by optimizing traditional Natural Language Processing (NLP) pipelines, improving performance and generalization. This has spurred their integration into various systems. Many NLP systems, including ours, employ a "one-stage" pipeline directly incorporating LLMs. While effective, this approach incurs substantial costs and latency due to the need for large model parameters to achieve satisfactory outcomes. This paper introduces a three-stage cost-efficient end-to-end LLM deployment pipeline—including prototyping, knowledge transfer, and model compression—to tackle the cost-performance dilemma in LLM-based frameworks. Our approach yields a super tiny model optimized for cost and performance in online systems, simplifying the system architecture. Initially, by transforming complex tasks into a function call-based LLM-driven pipeline, an optimal performance prototype system is constructed to produce high-quality data as a teacher model. The second stage combine techniques like rejection fine-tuning, reinforcement learning and knowledge distillation to transfer knowledge to a smaller 0.5B student model, delivering effective performance at minimal cost. The final stage applies quantization and pruning to extremely compress model to 0.4B, achieving ultra-low latency and cost. The framework’s modular design and cross-domain capabilities suggest potential applicability in other NLP areas.
[…]
1 Introduction
In recent years, Large Language Models (LLMs) have made significant strides [1; 2; 3] in enhancing artificial intelligence by addressing shortcomings of traditional pipelines — including manual annotation, suboptimal performance and limited generalization and adaptability. Consequently, LLM-based pipelines have stimulated substantial demand for integration into various Natural Language Processing (NLP) systems. We find that numerous NLP systems, including our own, have directly adopted LLM into the pipeline, often referred to as the "one stage" pipeline. Despite the breakthroughs facilitated by this pipeline, it also incurs substantial cost increases [4] and high inference latency [5]. Within our NLP system, cost-performance is an especially critical metric, a sentiment likely shared by others in the field [6]. These significant costs limit the broader applicability of LLMs and impact their commercial viability. Our analysis points that it is precisely the simplistic training method of the one-stage pipeline, which necessitates larger model parameters to achieve satisfactory results, that leads to significant costs.
Figure 1: From the traditional BERT-based system to the LLM-based system.
In this paper, we introduces three-stage performant and cost-efficient end-to-end LLM deployment pipeline, consisting of prototyping, knowledge transfer, and model compression, to address the pervasive "cost-performance" dilemma in practical deployment of LLM-based frameworks. This approach ultimately yields a super tiny model, which is deployed in our online system, achieving optimal cost-performance metrics while substantially simplifying system architecture. The first stage involves constructing the prototype system, where traditional complex tasks are converted into function call-based LLM-driven workflows [7] through prompt engineering, resulting in an optimal performance prototype compared to traditional pipeline. This prototype serves as a teacher model in the second stage, efficiently generating vast quantities of high-quality data. During the knowledge transfer phase, leveraging the optimal prototype from the first stage, we employ techniques including rejection fine-tuning (RFT), reinforcement learning (RL) [8; 9; 10], and multi-strategy knowledge distillation (KD) targeting the Qwen 2.5 series model [11], effectively transferring knowledge to a 0.5B student model. The derived student model is considerably smaller than the prototype model while achieves nearly comparable performance at minimal cost. In the third stage, we utilize quantization and pruning techniques to further compress the student model developed in the second phase, resulting in a remarkably compact model of 0.4 billion parameters (as shown in Figure 1). Ultimately, this super tiny model is deployed in our online system. This model achieves hundreds-fold compression and extremely low latency and cost in exchange for a very small and acceptable loss in performance. The modular design and superior cross-domain performance of the framework suggest its potential applicability in other NLP domains. The major contributions of this paper are as follows:
  • The first validated three-stage cost-efficient end-to-end LLM deployment pipeline that resolves the "cost-performance" dilemma through integrated technical practice, offering new insights for potentially cross-domain industrial-level solutions.
  • Introduction of an LLM-based generation paradigm replacing BERT classification for sample generation and evaluation, that leverages the benefits of LLM technology in line with current trends and enhances NLP performance across multi-domain scenarios.
  • Extensive experiments demonstrating state-of-the-art results, achieving an 180-fold compression with nearly consistent AR compared to 72B LLMs and improving accuracy by an absolute increase of 13% compared to traditional BERT-based pipelines.
2 Related Works
LLM for Function Calling & Evaluation
Leveraging in-context learning and reasoning capabilities, LLMs can select functions and parameters from provided lists based on user intent [12]. Fine-tuned 7B models like Gorilla [13] achieve GPT-4-level Function Call accuracy on datasets with thousands of APIs. Building on this, TinyAgent [14] enables function calling capabilities for edge devices with sub-1.1B parameter models. Query-to-API reformulation via rejection sampling can also optimize API retrieval, reducing retrieval latency while improving accuracy [15]. Nowadays, LLMs can be used to evaluate model-generated content [16; 17]. CriticGPT [18] indicates that LLMs can build a cost-effective and high-precision judge, surpassing humans in detecting subtle errors.
Knowledge Distillation
Knowledge distillation (KD) [19] transfers knowledge from a teacher model to a smaller student model by simulating probability distribution. Common approaches include Logits-based[20; 21], Feature-based [22] and sequence-level (Seq-KD) [23] distillation. Initially, Logits-based KD uses the forward KL divergence (FKL) [24; 25] as the loss function. Some KD methods prefer Reverse KL divergence (RKL) [26; 27; 28; 21]. Recently, Adaptive Kullback-Leiber (AKL) divergence [29] achieves a balance between mimicking distribution and avoiding overfitting.
LLM Pruning & Quantization
Efficient LLM deployment primarily relies on pruning[30; 31] and quantization[32; 33; 34]. Structured pruning[35] employs layer removal[36; 37] or dimensionality reduction[31; 37; 38] using gradient-based[31; 38] or activation-aware[37] importance criteria, implemented via learnable masks[38; 39] or pruning-retraining pipelines[36; 31; 37]. Weight-only LLM quantization methods[32; 33; 40] improve memory efficiency while joint weight-activation schemes[34; 41] and their combination [42] accelerate computation.
Reinforcement Learning
Recent works [8; 9; 10] focus on leveraging Reinforcement Learning (RL) to enhance model reasoning capabilities. Notable reasoning models such as DeepSeek-R1 [43], QWQ-32B [44], and OpenAI o1 [45] have emerged, employing RL algorithms like Proximal Policy Optimization (PPO) [46], Group Relative Policy Optimization (GRPO) [47], and others [48; 49; 50]. For general-purpose LLMs, preference optimization algorithms [51; 52; 53; 54] are widely adopted.
3 Methodology
Figure 2: Overview of Prototype, Production and Application.
Recent advancements in LLMs have facilitated significant progress in NLP tasks such as classification and summarization. We harness these capabilities to reconstruct the traditional dialogue system through a cost-efficient end-to-end LLM pipeline, which includes prototyping, knowledge transfer, and model compression, as illustrated in 2(a) and 2(b). Initially, leveraging the exceptional performance of contemporary LLMs, we develop the prototype system. Subsequently, we perform knowledge transfer from large models to smaller models. Finally, we achieve further reduction in model size through additional model compression techniques. Ultimately, our objective is to deliver intelligent user experiences at minimal costs, as demonstrated in 2(c).
3.1 Prototyping
Table 1: Complexity statistics of Our dataset vs Open Source datasets. More details in Appendix A.
Dataset Test Set Size Intents Slots TTR Entropy
TinyAgent[14] 1023 17 48 0.0087 7.8220
Snips[55] 893 7 72 0.0207 7.6997
ATIS[56] 700 21 120 0.0111 7.4148
Our dataset 11922 74 225 0.1410 9.4250
LLM Prototype
Traditional dialogue systems typically employ the Intent Classification and Slot Filling (ICSF) framework, often with BERT[57]. However, a single model often fails to handle complex real-world issues, as shown in the Table 1. Therefore, a Domain Classifier (DC) is widely used to first categorize queries into domains, enabling domain-specific ICSF parsing. While the hierarchical DC-ICSF architecture improves accuracy, system complexity rises with feature updates. This results in high inter-module coupling, increasing development and maintenance costs. Moreover, ambiguous domain boundaries make the DC a bottleneck for end-to-end accuracy. This creates a "feature expansion - operational difficulty" feedback loop, hindering dialogue experience optimization.
Modern LLMs demonstrate sufficient capability to address such NLP problems. Particularly in related Tool Use works, LLMs can learn to call functions without training by attaching function descriptions in prompts. In fact, ICSF can essentially be viewed as a function call task, where the model extracts critical information from user queries as function parameters for specific modules to execute. Following this approach, we design a prototype system based on Qwen2.5-72B-Instruct[11]. As illustrated in 2(a), the prototype system accepts user queries and contextual information (including chat history and temporal status information) as input, searches relevant APIs from the API pool, then outputs function names and parameters to call. The traditional system comprises one DC module and multiple ICSF models (based on BERT or regular expressions), each requiring separate maintenance and frequent DC updates. In contrast, the LLM-based system with function call capability handles all processing through a unified model, as demonstrated in Figure 1, which reduces system complexity and improves operational efficiency.
To enhance temporal awareness and mitigate hallucinations, we enrich the input with contextual knowledge via RAG [58], retrieving explanations and extensions of entity-related terms from a curated knowledge base. Instead of adding all callable APIs in the API pool into the LLM prompt, we implement an API retrieval module inspired by Kachuee et al. [15]. The LLM first generates potential API descriptions, then retrieves top-K APIs via vector similarity (using GTE-Qwen2-1.5B [59] as the embedding model and FAISS [60] for indexing). The LLM selects the optimal function and parameters from candidate APIs. If no suitable function exists, the system returns "No appropriate function"; ambiguous queries trigger "Request unclear" responses. For backward compatibility, we design an Adapter to map functions/parameters to legacy domain-intent-slot structures.
Judge
The traditional evaluation scheme for ICSF relies on the test set that requires full alignment of Intent and Slots. However, in real-world scenarios, there exist multiple solutions for function and slot selection. For example: (1) "Slightly increase the volume" could be resolved using either volume setting or volume increasing APIs; (2) "11:00 pm" and "23:00" semantically represent the same time. Additionally, due to crowdsourced annotators’ limited understanding of specific business logic, the manual labeling accuracy remains suboptimal. As shown in 2(c), the crowdsourced labeler’s annotation accuracy has become the system bottleneck.
To address issues above, we introduce LLM-as-a-judge to evaluate prototype accuracy. We design a Judge based on QwQ-32B[44]: Given context and a set of Function & parameters, the Judge determines whether this combination semantically satisfies the user query. Following Samuylova’s best practices [^1] https://www.evidentlyai.com/llm-guide/llm-as-a-judge, we optimize the Judge using a 1k-expert-annotated dataset. Subsequently, both the LLM and 9 domain experts evaluate 3k test samples under identical conditions. Results show that Judge achieves 93% consistency with expert judgments, with LLM providing logically coherent Chain-of-Thought(CoT) rationales. Based on these findings, we conclude that LLM-as-a-Judge represents the optimal evaluation approach for this task. As demonstrated in Figure 4, we reconstruct the dialogue pipeline from BERT-based DC-ICSF architecture to LLM-based Function Call architecture. Evaluated through LLM-as-a-Judge, the new architecture improves accuracy from 81.43% to 94.74% compared to the legacy system.
3.2 Knowledge Transfer
Rejection sampling Fine-Tuning
Building on our agent-oriented large model design, we propose Supervised Fine-Tuning (SFT) on smaller models using large model-generated data instead of relying solely on in-context learning (ICL) with large models. This approach offers two key benefits: (1) Smaller models can achieve performance comparable to large models when trained on sufficient high-quality data; (2) It significantly boosts inference and optimization efficiency. For a dataset $D$ containing prompt component $x$ and corresponding response component $y$, we define $D(x,y)=\{x_{i},y_{i}\}_{i=1}^{N}$, where $N$ denotes the dataset size. To minimize the Cross Entropy (CE) loss, we first utilize the SFT optimization, and then employ further optimization Rejection sampling Fine-Tuning (RFT). We use LLM judge introduced in Section 3.1 to evaluate training data generated by the LLM prototype, filter out low-quality data, and generate the final training set. The actual training data is formulated as:
$D_{filter}=\{x_{i},y_{i}\lvert\ \text{judge\_data}(x_{i},y_{i})=1\}_{i=1}^{N}$ (1)
where
[…]
Reinforcement Learning
Since OpenAI introduced Reinforcement Learning from Human Feedback (RLHF) [61], it has become an essential component for aligning LLMs, as evidenced in major technical reports [62; 63; 64; 11; 65]. DeepSeek-R1 [43] proposed a cost-effective RL training method from scratch which combined rule-based ORM and GRPO [47]. Subsequent works like Logic-RL [8; 9; 10] also successfully applied this efficient approach across various tasks. We adapt this method to allow the model to automatically generate Chain of Thought (CoT) during training, which enhances performance [66] and provides insights into model reasoning for subsequent optimization. Also, we primarily follow Logic-RL [9], refining the reward design for this specific task (see Appendix B) while utilizing the same training dataset as RFT to achieve enhanced model performance.
Distillation
Using the Function Call-related data from large models and the RFT methods, we transfer capabilities from large to small models through training. To accelerate model training and inference, we implement prompt compression during training by removing unnecessary content from input-output pairs, including format specifications in inputs and Chain-of-Thought (CoT) rationales in outputs. Please refer to Appendix C for details.
Knowledge distillation [19] is an effective method for transferring knowledge from large to small models, requiring fewer data and exhibiting lower overfitting risks compared to SFT. We adopt the logits-based Adaptive Kullback-Leibler (AKL) distillation [29], with the optimization objective:
$L_{AKL}=\mathbb{E}_{(x,y)\sim D}[AKL(\pi_{\phi}(y\lvert x),\pi_{\theta}(y\lvert x))]$ (3)
$AKL(p,q)=\alpha_{head}KL(p,q)+(1-\alpha_{head})KL(q,p)$ (4)
$KL(p,q)=\sum_{k}{p_{i}}\log\frac{p_{i}}{q_{i}}$ (5)
where $\pi_{\theta}$ denotes the trainable student model, $\pi_{\phi}$ represents the teacher model, and $\alpha_{head}$ is the head ratio in AKL. Setting $\alpha_{head}$ = 1 yields Forward KL Divergence-based distillation [25], while $\alpha_{head}$ = 0 corresponds to Reverse KL Divergence-based distillation [21].
We make optimizations on original methods: (1) Teacher Calibration: Following [67], we calibrate the teacher model with task-specific data to achieve lower student model loss; (2) Logits Pre-storage: For reducing training time, we pre-store teacher logits to enable multi-student distillation from a single cached logits set; (3) Top-k Pruning: We trim low-probability logits through top-k pruning, mitigating storage pressure and accelerating training speed.
3.3 Model Compression
Pruning
To further compress model size, we design a two-stage pruning pipeline following NVIDIA [67], Llama3.2 [68], and Apple [69]. Offline One-Shot Struct Pruning: Following Minitron[31], we explore two pruning strategies, for Depth Pruning, we compute layer importance scores and remove the $N$ least important layers based on the target pruning ratio, for Width Pruning, we conduct a model architecture search using PPL-based criteria and prune channels according to group-wise importance rankings, with full implementation details in Section I.1. Recovery training: After offline one-shot pruning, we execute a two-phase recovery process: (1) General Instruction Fine-tuning via Knowledge Distillation: Adopting Minitron’s approach, we leverage logits from a larger general instruction teacher model for recovery training. Due to unavailability of the teacher’s original training data, we first perform model correction on our selected IFT dataset to stabilize distillation, followed by top-k logits distillation(Section 3.2); (2) Domain-Specific Task Alignment: The pruned compact LM from phase 1 undergoes further aligning through knowledge distillation, producing a final domain-optimized LLM.
Quantization
Additionally, quantization significantly reduces GPU memory consumption and enhances inference performance. In this work, we apply post-training quantization to the smaller models obtained through distillation/pruning. Specifically, we implement two quantization approaches: 4-bit weight-only GPTQ algorithm [33] and full 8-bit weight-activation FP8 quantization [41]. The 4-bit weight-only quantization [33] compresses linear weights from FP16 to per-group INT4 via integer quantization, which are dequantized to FP16 during computation. This method achieves notable inference acceleration in memory-bound scenarios with large model parameters and low GPU memory bandwidth. The FP8 quantization [41] quantizes both weights and activations to 8-bit floating-point format. Compared to 8-bit integer quantization [41], FP8 demonstrates superior precision and leverages NVIDIA’s FP8 tensor cores for effective inference acceleration [70].
4 Experiments
4.1 Experimental Setup
Datasets
We conduct experiments using a large-scale dataset developed by our team, leveraging years of experience and practice in the field. The dataset transforms a real system’s intent recognition and semantic slot framework into an API document specification, creating a function pool with 74 functional modules and 225 parameters. We sample 200,000 data points from a single business scenario as the training set, with corresponding labels synthesized based on the large language model (LLM) prototype outlined in Section 3.1. Additionally, we employ a combination of hierarchical and random sampling to generate 11,922 samples as an in-domain test set, and 7,867 samples from a different business scenario as an out-domain test set, where some callable functions are absent from the training set. Compared to open-source datasets, our custom dataset is more intricate and closely mirrors real-world scenarios (see Table 1).
Metrics
In accordance with the LLM-based Judge outlined in Section 3.1, we establish our evaluation metrics. Consider a dataset of size $N$, denoted as $D(x,y)=\{x_{i},y_{i}\}_{i=1}^{N}$, where $x$ signifies the prompt, including elements such as user queries, function lists, and dialogue histories, while $y$ represents the response containing the function and its parameters. The function judge_data, as rigorously defined in Equation 2, assesses whether the function and its parameter list can fulfill the specified user query. We define the Achievable Rate (AR) as follows:
[…]
Training
We select small models (0.5B-7B) from the Qwen2.5 model family [11] for RFT, perform RL on the 3B model, and pruning on the 0.5B model, obtaining Qwen2.5-0.4B-Width and Qwen2.5-0.4B-Depth, both achieved under a 20% pruning ratio. Then, we perform KD on the pruned models and 0.5B model with the 7B RFT-trained teacher. Finally, we apply model quantization using W8A16/W4A16/W8A8 strategies to the KD-trained 0.5B model, detailed in Appendix E. The training process is cost-efficient, utilizing only 8 NVIDIA H20 GPUs. Computing resources are listed in Appendix F.
Evaluation settings
We evaluate our models with the BERT-Based System, which consists of two models based on BERT-Large (330M)[71]. To evaluate the effectiveness of our method, we invite expert annotators alongside the LLM-as-a-judge-based metric AR (see Equation 6) for performance assessment. We also implement a retrieval module to select the top 10 functions from the function pool, aiming to accelerate both the training and inference phases, as detailed in Appendix D. In real-world scenarios, we deploy models on NVIDIA L20 GPU and perform effectiveness evaluation via TensorRT-LLM[70], with specific configuration details provided in Appendix H. During the comparison between the trained models and a BERT-based system, we adopt queries per second (QPS) as the primary metric for system throughput, response time (RT) for end-to-end latency, and concurrent clients (CC) to quantify high-concurrency performance.
4.2 Experimental Results
Table 2: Main results of the models trained on Qwen2.5.(See Appendix J for FLOPs analysis)
Model / System AR GFLOPs
Crowdsourced Labelers 77.85% -
BERT-Based System 81.43% 38.79
Qwen2.5-72B (ICL) 94.74% 5223.91
Qwen2.5-7B (RFT) 95.29% 572.47
Qwen2.5-3B (RFT) 95.30% 297.98
Qwen2.5-1.5B (RFT) 95.16% 171.19
Qwen2.5-0.5B (RFT) 94.66% 77.89
Qwen2.5-0.5B (KD) 94.85% 77.89
Qwen2.5-0.5B-W8A16 (KD) 94.76% -
Qwen2.5-0.5B-W4A16 (KD) 94.19% -
Qwen2.5-0.5B-W8A8 (KD) 94.83% -
Qwen2.5-0.4B-Width (KD) 94.17% 72.93
Qwen2.5-0.4B-Depth (KD) 94.20% 67.68
Figure 3: Cost and performance results. FLOPs reduced by 10x with nearly lossless AR.
Table 3: RFT vs RL on in-domain data and out-domain data.
Model In-
Domain
Out-
Domain
AR
Qwen2.5-72B (ICL) 94.74% 94.63%
Qwen2.5-0.5B (RFT) 94.66% 89.32%
Qwen2.5-3B (RFT) 95.30% 91.02%
Qwen2.5-3B (RL) 94.68% 92.95%
Figure 4: Comparison over different models by professional annotators
LLM Prototype Preferred by Both Human and LLM-as-a-Judge
As illustrated in Figure 4, annotations provided by professional annotators reveal that the LLM prototype (72B ICL) achieves a win+tie rate of 95.59% over the BERT-based system and 92.98% over the crowdsource labeler. This demonstrates the superior performance of the LLM prototype. Using the LLM-as-a-Judge for evaluation, we find that the LLM Prototype surpasses the Bert-Based System by 13.31% and crowdsourced annotations by 16.89%, as displayed in Table 2. This conclusion is roughly consistent with those provided by professional annotators, demonstrating the effectiveness of the LLM judge.
RFT and KD as Effective Knowledge Transfer Methods
As demonstrated in Table 2, models trained with RFT generally exhibit a higher AR than the LLM prototype. Specifically, both the 3B and 7B models display negligible differences in performance but each achieve a 0.5% improvement in AR compared to the LLM prototype. In contrast, the 0.5B model shows a 0.5% lower AR than the 1.5B model. Moreover, KD emerges as a valuable complement to RFT, enhancing the performance of the 0.5B model by an additional 0.19%. In terms of inference performance, the 0.5B model demonstrates only a slight 0.4% difference in AR compared to the 7B model, while achieving a significant 13.57x improvement in QPS and a substantial 91.3% reduction in latency, as illustrated in Table 4.
RL’s Superior Generalization Ability
According to Table 3, while RFT performs well with in-domain data, it demonstrates poorer results with out-of-domain data, showing a 3.61% AR decrease compared to the LLM prototype. Conversely, RL models present a smaller gap (1.68%) with the LLM prototype on out-of-domain data, indicating better generalization of RL training methodologies.
Tiny LMs as a Cost-Efficient Upgrade to BERT-based System
Table 2 and Table 4 illustrate that quantization introduces a negligible AR loss of less than 0.1% while significantly enhancing inference performance. In the optimal scenario, quantization results in a 1.29x increase in QPS and a 0.76x decrease in RT. On the other hand, pruning results in a slightly higher accuracy loss, approximately 0.2%, but offers a more consistent 20% improvement in performance. In terms of end-to-end inference performance, under the condition where the CC is 64, the BERT-based system achieves a QPS of 1110.88 and an RT of 55.84 ms. Despite the cost being approximately doubled compared to the BERT-based system, the AR is improved by an absolute increase of 13%.
Table 4: Inference Performance under different model compression techniques.
Model QPS
cc=1 cc=8 cc=16 cc=32 cc=64 cc=128
Qwen2.5-0.5B 39.11 166.71 248.62 348.87 459.12 517.13
Qwen2.5-0.5B-W8A16 44.11 169.86 251.21 346.17 445.84 521.41
Qwen2.5-0.5B-W4A16 49.83 175.29 256.05 338.16 453.46 519.22
Qwen2.5-0.5B-W8A8 40.40 158.15 256.32 382.13 558.53 668.35
Qwen2.5-0.4B-Width 43.65 167.65 260.58 372.59 502.82 581.23
Qwen2.5-0.4B-Depth 45.22 198.82 309.62 438.58 566.87 650.26
Qwen2.5-7B 4.43 14.77 17.71 19.04 19.79 20.06
Figure 5: Normalized RT ratio based on FP16 models
5 Analysis
Performance-Efficiency Trade-offs in Model Scaling
Our experimental analysis highlights a compelling trade-off between model size and computational efficiency. Although larger models typically demonstrate superior performance, the expected disparity in performance between large and smaller models is not as pronounced as anticipated. As shown in Figure 3, the 0.4B model achieves performance almost on par with the 7B model, efficiently requiring 11.8% FLOPs, while even reaches more than 17.4x QPS improvement and 92.9% RT reduction, with the 7B model’s RT being 882.552 ms under 16 CC. Although this model has a slight gap in cost compared to traditional BERT models, the attractive accuracy and scalability stemming from LLM base makes it a favorable choice for various applications.
Characteristics of Training Paradigm
The results presented in Table 3 underscore the distinct characteristics of different training methodologies. Fine-tuning primarily focuses on memorizing data, making it susceptible to overfitting to the training set distribution. In contrast, RL demonstrates greater robustness by not merely fitting individual data points. We posit that RL’s superior generalization capabilities stem from its ability to learn from negative samples, which are generated by the model being trained itself. RL is highly effective for dealing with well-structured problems, while it presents unsatisfactory convergence facing ambiguous data or a lack of domain knowledge in real-world tasks. Moreover, RL is sensitive to noisy data. Despite these limitations, RL remains a promising direction for development, especially when building upon the foundation of traditional fine-tuning training strategies.
Comparison over Model Compression Strategies
Our analysis identifies depth pruning as the optimal approach for domain-specific models. Depth pruning achieves 20% inference acceleration with lossless AR at equivalent compression ratios (Table 2), stemming from the nature of depth pruning, removing entire transformer layers and their associated attention computations. Although width pruning exhibits slightly better effects in general instruction tasks after training (Table 9), depth pruning demonstrates optimal efficiency facing long prompts, with width pruning focusing on matrix optimizations (Appendix J). Regarding quantization, the W4A16 strategy achieves impressive performance in single-thread scenarios, while FP8 outperforms under high concurrency conditions with lossless accuracy, considered a preferred deployment paradigm, highlighting the importance of throughput.
6 Conclusion
Based on our comprehensive study, several key insights emerge regarding the development and deployment of effective deep learning models in practical scenarios.
Tiny Models as a Practical Low-Cost Alternative
Our experiments indicate that end-to-end optimized tiny models can achieve performance levels comparable to their larger counterparts, offering resource efficiency without compromising accuracy. This provides an alternative pathway for applying LLMs in practical tasks under low-cost scenarios.
Balanced Focus Beyond AGI
While AGI remains a visionary long-term aim for LLMs, progress has reached a plateau, prompting increased attention towards real-world applicability. Our work demonstrates that integrating LLM-driven techniques into traditional tasks results in immediate, measurable improvements. It advocates for a balanced research agenda that equally emphasizes foundational advancements and domain-specific applications.
Cost-Efficient Three-Stage Pipeline
The proposed pipeline—comprising prototyping, knowledge transfer, and model compression—presents a systematic framework designed to achieve high performance while maintaining low costs in both training and inference. We believe that this approach can be effectively adapted to other domains, thereby potentially advancing research in those areas and facilitating the widespread deployment of LLMs.
7 Limitation & Future Works
While our proposed three-stage pipeline demonstrates promising results in cost-performance, several limitations warrant further investigation, underscoring the need for holistic optimization across algorithmic robustness, dynamic adaptation and hardware efficiency to enable truly scalable and trustworthy adoption in industrial settings:
Continual Learning for Dynamic Adaptation
While cost-efficient, the static training-deployment cycle of our pipeline hinders its adaptation to dynamic tasks and evolving datasets. Future work should develop advanced continual learning frameworks that enable incremental knowledge integration without catastrophic forgetting, leveraging rapid training cycles to maintain deployment agility. Employing online learning and lifelong learning will enhance models’ robustness and adaptability, ensuring sustained relevance and precision in changing linguistic contexts and shifting user demands.
Edge Deployment Optimization with Performance Assurance
Being sufficiently compact, our super tiny models have demonstrated impressive capabilities in currently high-demand applications of IoT devices, such as mobile phones and tablets, characterized by low latency, cost-effectiveness, cloud independence, and robust privacy. While further compression to an even smaller size like 0.1B may not be significantly beneficial, exploration of optimizing on-device deployment of such ultra-small models remains a promising direction for future research. Future works should develop hardware-aware compression algorithms and adaptive runtime kernels to optimize latency-energy tradeoffs across diverse IoT ecosystems, addressing real-world heterogeneous hardware constraints such as memory fragmentation and NPU compatibility. We will also explore more resource-efficient deployment techniques for edge devices—such as sparse computation, edge-centric quantization, and federated learning.
Appendix A Dataset Analysis
Type-Token Rate (TTR) is a fundamental measure of lexical diversity that reflects the variability of vocabulary locally, but it is sensitive to text length. The calculation formula is:
$TTR=\frac{N_{types}}{N_{tokens}}$ (7)
where $N_{tokens}$ denotes the total number of 2-gram patterns in the text, and $N_{types}$ represents the number of distinct 2-gram patterns.
Entropy is an information-theoretic metric that quantifies the uncertainty within binary grammatical sequences, based on their probability distribution. A higher entropy value indicates more complex local language patterns. The calculation formula is:
$Entropy=-\sum_{i}^{N_{types}}p_{i}\log_{2}(p_{i})$ (8)
where $p_{i}$ is the probability of occurrence for the $i^{th}$ 2-gram pattern.
Both TTR and Entropy are sensitive to text length. To enable a relatively fair comparison of the complexity across datasets, we randomly sample 700 entries from the test set of each dataset for the computation(as the ATIS dataset only contains 700 instances).
Appendix B Details for Reinforcement Learning
Our reward design consists of two components, Format Reward($S_{format}$) and Answer Reward ($S_{answer}$):
  • For Format Reward, we adopt the logic of Logic-RL[9], which utilizes regular expressions to validate the output format against the specified template <think>...</think><answer>...</answer>. Additionally, we extend this design by verifying the legality of function usage and its parameters. For example, if it choose a function that doesn’t exist in the input function pool, then it is considered as invalid.
$S_{format}=\begin{cases}1,&\text{if the format is correct, and the function call is valid}\\ 0,&\text{otherwise}\end{cases}$ (9)
  • For Answer Reward, according to our specific scenario, we find that the model attempts to directly guess answer during the training, leading no association between the Chain-of-Thought (CoT) and the final results. Requiring explicit repetition of the answer within the CoT framework effectively resolves this issue. As demonstrated in Figure 6, we optimize its original design in order to make the Chain-of-Thought (CoT) reasoning and the results more consistent. We require: (1) the CoT must explicitly reference the target function at least once; (2) the final function mentioned in the CoT must match the function in the answer. Otherwise, it will be considered as a parsing error.
$S_{answer}=\begin{cases}2,&\text{if the answer is correct,}\\ &\text{and the thought ends with the function in the answer,}\\ -2&\text{if the answer cannot be parsed or is missing,}\\ &\text{or is not defined in the function list,}\\ &\text{or the thought doesn't end with the function in the answer}\\ -1.5,&\text{otherwise}\end{cases}$ (10)
Figure 6: An example of optimization of the original Answer Reward design by requiring explicit repetition of the answer within the CoT framework.
Appendix C Prompt Compression
To reduce inference costs, we remove unnecessary content in prompts, such as guidelines, format specifications, and few-shot examples used during large model invocation, while retaining essential elements: user queries, context, query-relevant knowledge, function lists, and brief function descriptions. For outputs, we remove Chain-of-Thought (CoT) components and directly use predefined JSON-formatted responses, as illustrated in Figure 7.
Figure 7: Original vs Compressed Prompt & Response.
Appendix D Retriever Setup
Initially, we employ GTE-large [59] as the embedding model trained based on the BERT-large [71]. We compare the cosine similarity of the embedding between the API description generated by LLM based on this model and the API description in the API document, but the recall accuracy is not high.
Subsequently, we replace it with GTE-Qwen2-1.5B-instruct [59] trained based on Qwen2-1.5B [97]. This model has the advantage of being able to fine tune matching tasks by modifying prompts.
Next, based on this model and a customized prompt in Section D.1, we achieve a higher accuracy by matching the cosine similarity between the API description generated by the model and the JSON text of the entire API document. During training, we optimize the API description model through rejection sampling, with implementation details provided in Section D.2.
Finally, to achieve maximal inference speed, we trained a classifier based on BERT-large [71] for top-1 recall using rejection sampling. The classifier takes Dialog, User Query, and necessary context as input and outputs the function category. Then, the functions corresponding to the top-k output logits are considered as candidate functions for the subsequent model’s function call. This approach significantly improves inference speed (due to the inherent efficiency of BERT classifiers compared to autoregressive GPT models) and also enhances function recall. However, this comes at the cost of scalability, as retraining the recall model is required for each update to the function pool.
As illustrated in Table 5, these optimizations substantially improve API recall rate.
Table 5: Comparison of retrieval performance of models: GTE-large model, GTE-Qwen2-1.5B-instruct model, optimized API description model + GTE-Qwen2-1.5B-instruct model. Extensibility implies that the API pool can be dynamically modified.
Top-N GTE-large GTE-Qwen2 Opt-API+GTE-Qwen2 RFT-BERT
5 60.19% 88.27% 93.39% 99.61%
7 74.06% 94.67% 98.09% 99.92%
10 90.11% 98.22% 99.55% 99.92%
15 98.95% 99.80% 100% 100%
Extensibility ×
D.1 Prompt for retrieval
{mdframed}
Instruction: Given a API description, retrieve the most similar API from the function pool
Query:
D.2 Retriever Optimization
Algorithm 1 Simplified API Description Validation Process
0:  Query $q$, LLM $\mathcal{M}$, API Pool $\mathcal{R}$, Retrieval Module $\mathcal{S}$, Function Call Module $\mathcal{F}$, Judge $\mathcal{J}$
0:  High-quality training dataset $\mathcal{D}$
1:  Generate multiple API descriptions $Des=\{\mathcal{M}(q)\}$ in parallel.
2:  for each description $d\in Des$ do
3:     Retrieve candidate APIs $C=\mathcal{S}(d,\mathcal{R})$.
4:     Select API $a\in C$ and generate parameters $p$.
5:     Execute $a$ with $p$, get result $r$.
6:     if $\mathcal{J}(r)=\text{Success}$ then
7:        Check if $a$ is in the top 5 candidates of $C$.
8:        if Yes then
9:           Add $(q,d,a,p,r)$ to $\mathcal{D}$.
10:        end if
11:     end if
12:  end for
Appendix E Training Details
Table 6: Training Parameters in different stages.
Supervised Fine-tuning (SFT)
Learning Rate $2\times 10^{-5}$
Batch Size $128$
Number of Epochs $1$
Weight Decay $1\times 10^{-4}$
Reinforcement Learning (RL)
Learning Rate $3\times 10^{-7}$
Batch Size $128$
Number of Episodes $1400$
Number of Rollouts $8$
KL Coefficient $0.001$
Knowledge Distillation (KD)
Learning Rate $2\times 10^{-5}$
Batch Size $128$
Number of Epochs $1$
Logits Top-K $100$
Table 7: Different KD results on Qwen2.5-0.5B model.
Model AR
Qwen2.5-0.5B (RFT) 94.66%
Qwen2.5-0.5B (FKL) 94.65%
[…]
The main training parameters are shown in Table 7. During the training phase, we utilize the AdamW optimizer [96], setting the model’s maximum sequence length to $5120$. In the RL phase, we employ the baseline variant of Reinforce++ [50] as the reinforcement learning algorithm, offering advantages of efficiency, conciseness, and training stability. Following [121], the K2 estimator is used instead of the original K1 in Reinforce++. In the AKL phase, the threshold parameter $\mu$ is reduced from 1 to 0.5. For pruning, we use the 0.5B-Instruct model as the base model and employ width/depth pruning strategies with 20%/30% prune ratio, finally obtaining two models Qwen2.5-0.4B(0.35B)-Width and Qwen2.5-0.4B(0.35B)-Depth. We produce the extremely compressed models through a two-stage training process (accuracy recovery training, KD). Complete training details for the pruned models can be found in Appendix I. For quantization, we primarily experiment with three quantization strategies on the 0.5B model trained with AKL: W8A16, W4A16, and W8A8. Both W8A16 and W4A16 adopt per-group symmetric int8 weight-only quantization with a group_size=128, and activations are in fp16. To improve accuracy, W4A16 use 512 training data samples combined with the GPTQ algorithm [33] for calibration. In contrast, W8A8 applies FP8 for both weights and activations. For the training frameworks, we use Megatron-LM [73] for SFT and KD, while OpenRLHF [74] is employed for RL. Data sampling during training is implemented using vLLM [72]. In the evaluation phase, we utilize SGLang [75] to generate data.
Appendix F Computing resources
We utilize 8 NVIDIA H20 GPUs for training process and data generation. For inference performance evaluation, we employ a single NVIDIA L20 GPU. We apply an INTEL XEON PLATINUM 8575C processor with 192 cores for computation, which operates at a frequency of 3028.956 MHz and a cache size of 327680 KB. In addition, we establish a virtual development environment based on Kubernetes, leveraging 64 virtual cores and 512 GB of memory.
During the training phase, the RFT process takes approximately 2 to 14 hours for models ranging from 0.5B to 7B. Pre-KD logits caching for a 7B model requires about 5 hours, while RL for the 3B model takes approximately 3 days. The KD process for the 0.5B model is completed in about 2 hours. Training each pruned model demands roughly 3.5 days.
Furthermore, for the LLM prototype and LLM-as-a-Judge, data generation of every 10,000 entries requires approximately 1 hour.
Appendix G Detailed Analysis for Experimental Results for KD
As demonstrated in Table 7, it is evident that balancing FKL and RKL is quite crucial for our task. The process of function selection by the model resembles a classification task, whereas the generation of parameter names and values for function calls is more akin to a generation task. The parameter $\mu$ in AKL is precisely used to adjust the effects between classification and generation. Experimental results demonstrate that at $\mu\leq 0.7$ and $\mu=1.0$ (equivalent to FKL), the performance with AKL is similar to or worse than RFT. Based on our findings where the average probability of the top-1 token from the teacher model exceeds 99%, we draw the conclusion that setting a lower $\mu$ value causes the effect of AKL to excessively shift more towards RKL. As described in the AKL paper, RKL prioritizes learning the tail of the probability distribution, which is detrimental for rapid convergence in scenarios without requiring diversity. However, RKL can bring considerable benefits for the function classification part. Overall, AKL demonstrates greater potential compared to RFT, achieving some performance improvements with an appropriately chosen $\mu$ value.
Appendix H Inference Details
The inference framework utilizes TensorRT-LLM[70] v0.16.0 with enabled optimizations including paged attention, prefix cache, CUDA graph size set to 1024, and max token number set to 32768 for inference acceleration.
Appendix I Prune Detail
I.1 Offline One-shot Pruning
Importance Metrics for Structural Pruning
(1) Depth Pruning: we propose Layer Importance (LI) metric based on the similarity between consecutive layers’ activation. Formally, we define the importance of the $i$-th layer as:
$LI_{i}=1-E_{X,t}\frac{X_{i,t}^{T}X_{i+1,t}}{\|X_{i,t}\|_{2}\|X_{i+1,t}\|_{2}}$ (11)
where $X_{i}$ denotes the activation matrix of layer $i$, with $X_{i,t}$ representing the $t$-th column (embedding vector) of $X_{i}$. After we collecte all layer activations $\{X_{i}\}_{i=1}^{L}$ through a single forward pass on calibration data, then remove the $N$ least important layers, where $N$ is determined by the target pruning ratio; (2) Width Pruning: We categorize width pruning along three dimensions by channel dependencies: Embedding dimension, Head dimension and Intermediate dimension(as as illustrated in ). For compact LLMs (e.g., Qwen2.5-0.5B with head size 14), we exclude head dimension pruning due to its demonstrated sensitivity[37], we use an activation-based importance metric Channel Importance(CI) to do Channel selection as:
$CI^{i}_{\text{embedding}}=\sqrt{\sum_{B,S}|\text{LN}(X)^{i}|^{2}},\quad CI^{i}_{\text{inter}}=\sqrt{\sum_{B,S}|x(W_{ffn1}^{i})^{T}|^{2}}$ (12)
where $CI^{i}_{\text{embedding}}$ denotes the importance score of the $i$-th embedding dimension, computed from LayerNorm outputs of attention and FFN blocks, while $CI^{i}_{\text{inter}}$ represents the importance of the $i$-th intermediate dimension, calculated from the $i$-th row of FFN’s gate and up projection matrices $W_{ffn1}^{i}$. The importance scores are first aggregated across Batch and Sequence dimensions through summation $\sum_{B,S}$, then normalized using L2 normalization within each subgroup to obtain global channel importance.
Figure 8: Pruning dependency in Qwen
Best Width Pruning Model Configuration
To maintain hardware-efficient matrix multiplication during inference, we align pruning granularity to 64 for Embedding dimension and 256 for intermediate dimension, yielding multiple width-pruned model configurations as shown in Table 8. We compute channel importance scores $CI^{(i)}_{\text{embedding}}$ and $CI^{(j)}_{\text{inter}}$ on a calibration dataset comprising 1024 randomly sampled sequences from WikiText-2. For each configuration, we prune entire channel groups with the lowest CI scores and evaluate the resulting one-shot pruned models through WikiText-2 perplexity. Table 8 presents the optimal configurations for width pruning at rates of 20% and 30%, which are identified as Qwen2.5-0.4B-Width-2 and Qwen2.5-0.35B-Width-3, respectively.
Table 8: Perplexity (PPL) on WikiText-2 for different model architectures with 20% and 30% width pruning.
Hidden
Size
FFN
Inter Size
Non-Emb
Params(B)
Emb
Param(B)
Params
WikiText2
PPL
Qwen2.5-0.5B 896 4864 0.358 0.136 0.345 14.28
Qwen2.5-0.4B-Width-1 896 3328 0.259 0.136 0.395 349.89
Qwen2.5-0.4B-Width-2 832 3840 0.271 0.126 0.397 251.95
Qwen2.5-0.4B-Width-3 768 4352 0.278 0.117 0.395 354.74
Qwen2.5-0.4B-Width-4 706 4864 0.282 0.107 0.389 493.55
Qwen2.5-0.35B-Width-1 896 2560 0.209 0.136 0.345 1210.42
Qwen2.5-0.35B-Width-2 768 3584 0.236 0.117 0.351 1141.69
Qwen2.5-0.35B-Width-3 768 3584 0.236 0.117 0.353 1009.57
Qwen2.5-0.35B-Width-4 706 4096 0.242 0.107 0.349 2026.00
Qwen2.5-0.35B-Width-5 640 4864 0.256 0.097 0.353 2395.52
I.2 Recovery Training in Pruning
General Instruction Fine-tuning Dataset And Traing Hyperparameters
We employ the Infinity-Instruct-7M[106] dataset and adopt AKL distillation[29] described in Section 3.2 for general instruction recovery training. We utilize Qwen2.5-7B-Instruct[11] as the teacher model, following Minitron’s methodology[67] to perform teacher correction on the Infinity-Instruct-7M dataset. We employ the Adam optimizer (weight decay=1e-4) with a cosine learning rate schedule decaying from 2e-5 to 2e-6. We conduct training for 2 epochs, utilizing a batch size of 32, and each batch is packed to accommodate sequences of up to 7168 tokens.
Table 9: Performance of different pruned model compared to the unpruned model. We report 5-shot performance on mmlu and cmmlu, 4-shot for gsm8k, zero-shot for ARC-Challenge, HellaSwag. * indicates that the baseline Qwen-2.5-0.5B utilizes the same instruct fine-tuning dataset for correction.
model
MMLU
5-shot
CMMLU
5-shot
Arc-C
0-shot
HellaSwag
0-shot
GSM8k
4-shot
Avg.
Qwen-2.5-0.5B* 44.65 42.53 50.17 39.13 45.11 44.32
Qwen-2.5-0.4B-Width 38.63 35.89 42.27 35.55 48.60 40.12
Qwen-2.5-0.35B-Width 35.85 33.18 40.21 32.21 44.35 37.16
Qwen-2.5-0.4B-Depth 32.89 32.41 37.11 33.69 37.68 34.76
Qwen-2.5-0.35B-Depth 25.46 24.42 26.12 25.21 5.61 21.36
Evaluation on Knowledge & Reasoning Tasks
Table 9 presents the performance of pruned models on downstream tasks, including MMLU and CMMLU for knowledge assessment, as well as ARC-Challenge, HellaSwag and Winogrande for commonsense reasoning evaluation. Our experiments demonstrate distinct recovery patterns between width and depth pruning on general tasks. At 20% pruning rate, width pruning achieves 90.5% performance retention compared to the unpruned baseline, while depth pruning retains only 78.4%. When increasing pruning intensity to 30%, width pruning maintains 83.8% of unpruned performance, whereas depth pruning suffers catastrophic performance degradation-notably on GSM8K, where accuracy drops from 45.11 to merely 5.61.
Ablation of Pruning Ratio on Domain-Specific Task
Our investigation reveals a nonlinear relationship between pruning intensity and model accuracy in domain-specific tasks. Table 10 demonstrates that both depth-wise and width-wise pruning with 20% pruning rate incur negligible AR degradation (<0.5%). However, when increasing the pruning rate to 30%, we observe significant performance deterioration (>1.5% AR drop). These findings suggest that for small-scale models operating on domain-specific tasks, pruning rates exceeding 20% may lead to substantial accuracy compromises.
Table 10: Results of different pruning ratios and pruning strategies on domain specific task.
Model AR
Qwen2.5-0.5B (KD) 94.85%
Qwen2.5-0.4B-Width (KD) 94.17%
Qwen2.5-0.35B-Width (KD) 93.71%
Qwen2.5-0.4B-Depth (KD) 94.20%
Qwen2.5-0.35B-Depth (KD) 93.75%
Appendix J FLOPs Calculation
For the calculation of FLOPs, we first define the notations symbolically:
  • $B$: batch size
  • $H$: hidden size
  • $S$: sequence length (average number of prompt tokens)
  • $S_{\text{uncached}}$: average number of tokens in prefill phase not hit by prefill cache (requires recomputing $kv_{\text{cache}}$)
  • attn_heads: number of attention heads
  • kv_heads: number of separate key/value heads in GQA
  • head_size: dimension of each head
  • inter_size: intermediate dimension of FFN layer
  • vocab_size: vocabulary size
  • $L$: number of model layers
Following [122], we consider only the floating-point operations (FLOPs) from matrix multiplications (GEMMs). For the Qwen model:
  • Attention block operations:
    • Query/Key/Value transformations (GQA):
[…]
    • Attention matrix computation and attention over values:
[…]
    • Post-attention linear projection:
[…]
  • Feed-forward network block operations:
    • Up/Gate/Down projections (each):
[…]
Thus, the total FLOPs for one inference with prefill cache is:
[…]
Workload and Computational Analysis
Our experiments workload consists of prompts with an average length of 1792 tokens. By leveraging prefix caching to exploit shared prompt prefixes during inference, we reduce the average number of uncached tokens in the prefill phase to 128, while generating an average of 9 tokens per request. For FLOPs calculations, we account for $S_{\text{uncached}}=128$ tokens in the prefill phase and $S_{\text{uncached}}=1$ token per decoding step, with detailed results presented in Table 2.

### v2

@@
 While effective, this approach incurs substantial costs and latency due to the need for large model parameters to achieve satisfactory outcomes.
-This paper introduces a three-stage cost-efficient end-to-end LLM deployment pipeline—including prototyping, knowledge transfer, and model compression—to tackle the cost-performance dilemma in LLM-based frameworks.
+This paper introduces a three-stage cost-efficient end-to-end LLM deployment pipeline—including prototyping, knowledge transfer, and model compression-to tackle the cost-performance dilemma in LLM-based frameworks.
 Our approach yields a super tiny model optimized for cost and performance in online systems, simplifying the system architecture.
 Initially, by transforming complex tasks into a function call-based LLM-driven pipeline, an optimal performance prototype system is constructed to produce high-quality data as a teacher model.
-The second stage combine techniques like rejection fine-tuning, reinforcement learning and knowledge distillation to transfer knowledge to a smaller 0.5B student model, delivering effective performance at minimal cost.
-The final stage applies quantization and pruning to extremely compress model to 0.4B, achieving ultra-low latency and cost.
+The second stage combines techniques like rejection fine-tuning, reinforcement learning, and knowledge distillation to transfer knowledge to a smaller 0.5B student model, delivering effective performance at minimal cost.
+The final stage applies quantization and pruning to extremely compress models to 0.4B, achieving ultra-low latency and cost.
 The framework’s modular design and cross-domain capabilities suggest potential applicability in other NLP areas.
-[…]
-[…]
+* Equal Contribution Corresponding Author
 1 Introduction
-In recent years, Large Language Models (LLMs) have made significant strides [1; 2; 3] in enhancing artificial intelligence by addressing shortcomings of traditional pipelines — including manual annotation, suboptimal performance and limited generalization and adaptability.
+Figure 1: From the traditional BERT-based system to the LLM-based system.
+In recent years, Large Language Models (LLMs) have made significant strides [1; 2; 3] in enhancing artificial intelligence by addressing shortcomings of traditional pipelines — including manual annotation, suboptimal performance, and limited generalization and adaptability.
 Consequently, LLM-based pipelines have stimulated substantial demand for integration into various Natural Language Processing (NLP) systems.
@@
 These significant costs limit the broader applicability of LLMs and impact their commercial viability.
-Our analysis points that it is precisely the simplistic training method of the one-stage pipeline, which necessitates larger model parameters to achieve satisfactory results, that leads to significant costs.
-Figure 1: From the traditional BERT-based system to the LLM-based system.
-In this paper, we introduces three-stage performant and cost-efficient end-to-end LLM deployment pipeline, consisting of prototyping, knowledge transfer, and model compression, to address the pervasive "cost-performance" dilemma in practical deployment of LLM-based frameworks.
+Our analysis points out that it is precisely the simplistic training method of the one-stage pipeline, which necessitates larger model parameters to achieve satisfactory results, that leads to significant costs.
+In this paper, we introduce a three-stage performant and cost-efficient end-to-end LLM deployment pipeline, consisting of prototyping, knowledge transfer, and model compression, to address the pervasive "cost-performance" dilemma in the practical deployment of LLM-based frameworks.
 This approach ultimately yields a super tiny model, which is deployed in our online system, achieving optimal cost-performance metrics while substantially simplifying system architecture.
@@
 During the knowledge transfer phase, leveraging the optimal prototype from the first stage, we employ techniques including rejection fine-tuning (RFT), reinforcement learning (RL) [8; 9; 10], and multi-strategy knowledge distillation (KD) targeting the Qwen 2.5 series model [11], effectively transferring knowledge to a 0.5B student model.
-The derived student model is considerably smaller than the prototype model while achieves nearly comparable performance at minimal cost.
+The derived student model is considerably smaller than the prototype model while achieving nearly comparable performance at minimal cost.
 In the third stage, we utilize quantization and pruning techniques to further compress the student model developed in the second phase, resulting in a remarkably compact model of 0.4 billion parameters (as shown in Figure 1).
 Ultimately, this super tiny model is deployed in our online system.
-This model achieves hundreds-fold compression and extremely low latency and cost in exchange for a very small and acceptable loss in performance.
+This model achieves hundreds-fold compression and extremely low latency and cost in exchange for a trivial and acceptable loss in performance.
 The modular design and superior cross-domain performance of the framework suggest its potential applicability in other NLP domains.
 The major contributions of this paper are as follows:
-  • The first validated three-stage cost-efficient end-to-end LLM deployment pipeline that resolves the "cost-performance" dilemma through integrated technical practice, offering new insights for potentially cross-domain industrial-level solutions.
-  • Introduction of an LLM-based generation paradigm replacing BERT classification for sample generation and evaluation, that leverages the benefits of LLM technology in line with current trends and enhances NLP performance across multi-domain scenarios.
-  • Extensive experiments demonstrating state-of-the-art results, achieving an 180-fold compression with nearly consistent AR compared to 72B LLMs and improving accuracy by an absolute increase of 13% compared to traditional BERT-based pipelines.
+  • The first validated three-stage cost-efficient end-to-end LLM pipeline with simplifying original system complexity achieves SOTA results and resolves the "cost-performance" dilemma, realizing a 180-fold compression with nearly consistent AR compared to 72B LLMs and up to an absolute 14% AR increase compared to traditional BERT-based pipelines.
+  • We propose the first hybrid strategy that significantly enhances the generalization of super tiny models on domain-specific tasks through combining RL and KD.
+This approach effectively transfers the capabilities of LLMs to smaller models.
+  • The proposed framework offers advantages such as low training, inference, and deployment costs, minimal project development expenses with fast cycles, and a strong potential for cross-domain transferability.
+It provides inspiration for developing standardized cross-domain solutions in both academia and industry, facilitating the widespread LLMs adoption.
 2 Related Works
@@
 For general-purpose LLMs, preference optimization algorithms [51; 52; 53; 54] are widely adopted.
+Additionally, self-improving pipelines [55; 56; 57] have been proposed to iteratively refine model performance.
+We primarily utilize RL to enhance the model’s generalization.
+Compared to directly distilling a reasoning model, RL results in shorter and more precise Chains-of-Thought (CoT).
 3 Methodology
@@
 We harness these capabilities to reconstruct the traditional dialogue system through a cost-efficient end-to-end LLM pipeline, which includes prototyping, knowledge transfer, and model compression, as illustrated in 2(a) and 2(b).
-Initially, leveraging the exceptional performance of contemporary LLMs, we develop the prototype system.
+Initially, we develop the prototype system by leveraging the exceptional performance of contemporary LLMs.
 Subsequently, we perform knowledge transfer from large models to smaller models.
-Finally, we achieve further reduction in model size through additional model compression techniques.
+Finally, we achieve additional reduction in model size through further model compression techniques.
 Ultimately, our objective is to deliver intelligent user experiences at minimal costs, as demonstrated in 2(c).
@@
 LLM Prototype
-Traditional dialogue systems typically employ the Intent Classification and Slot Filling (ICSF) framework, often with BERT[57].
-However, a single model often fails to handle complex real-world issues, as shown in the Table 1.
-Therefore, a Domain Classifier (DC) is widely used to first categorize queries into domains, enabling domain-specific ICSF parsing.
-While the hierarchical DC-ICSF architecture improves accuracy, system complexity rises with feature updates.
-This results in high inter-module coupling, increasing development and maintenance costs.
-Moreover, ambiguous domain boundaries make the DC a bottleneck for end-to-end accuracy.
-This creates a "feature expansion - operational difficulty" feedback loop, hindering dialogue experience optimization.
-Modern LLMs demonstrate sufficient capability to address such NLP problems.
-Particularly in related Tool Use works, LLMs can learn to call functions without training by attaching function descriptions in prompts.
-In fact, ICSF can essentially be viewed as a function call task, where the model extracts critical information from user queries as function parameters for specific modules to execute.
+Traditional dialogue systems usually employ the Intent Classification and Slot Filling (ICSF) framework, often utilizing BERT[60; 61].
+However, a single model often struggles to handle complex real-world issues, as shown in the Table 1.
+Consequently, a Domain Classifier (DC) is frequently used to first categorize queries into domains, enabling domain-specific ICSF parsing.
+Although the hierarchical DC-ICSF architecture improves accuracy, the system complexity increases with feature updates.
+This leads to high inter-module coupling, increasing development and maintenance costs.
+Furthermore, ambiguous domain boundaries turn the DC into a bottleneck for overall end-to-end accuracy, creating a "feature expansion - operational difficulty" feedback loop that hinders dialogue experience optimization.
+Modern LLMs exhibit adequate capability to tackle these NLP problems.
+Particularly in related Tool Use works, LLMs can learn to call functions without additional training by attaching function descriptions in prompts.
+Essentially, ICSF can be viewed as a function call task, where the model extracts critical information from user queries as function parameters for specific modules to execute.
 Following this approach, we design a prototype system based on Qwen2.5-72B-Instruct[11].
@@
 The traditional system comprises one DC module and multiple ICSF models (based on BERT or regular expressions), each requiring separate maintenance and frequent DC updates.
-In contrast, the LLM-based system with function call capability handles all processing through a unified model, as demonstrated in Figure 1, which reduces system complexity and improves operational efficiency.
+In contrast, the LLM-based system, equipped with function call capability, processes all tasks through a unified model, as illustrated in Figure 1, thereby reducing system complexity and enhancing operational efficiency.
 To enhance temporal awareness and mitigate hallucinations, we enrich the input with contextual knowledge via RAG [62], retrieving explanations and extensions of entity-related terms from a curated knowledge base.
-Instead of adding all callable APIs in the API pool into the LLM prompt, we implement an API retrieval module inspired by Kachuee et al.
-[15].
+Instead of adding all callable APIs from the API pool into the LLM prompt, we implement an API retrieval module inspired by [15].
 The LLM first generates potential API descriptions, then retrieves top-K APIs via vector similarity (using GTE-Qwen2-1.5B [63] as the embedding model and FAISS [64] for indexing).
-The LLM selects the optimal function and parameters from candidate APIs.
-If no suitable function exists, the system returns "No appropriate function"; ambiguous queries trigger "Request unclear" responses.
+Then, the LLM selects optimal functions/parameters from candidate APIs.
+If no suitable function exists, the system returns "No appropriate function"; in the case of ambiguous queries, "Request unclear" are generated.
 For backward compatibility, we design an Adapter to map functions/parameters to legacy domain-intent-slot structures.
@@
 Subsequently, both the LLM and 9 domain experts evaluate 3k test samples under identical conditions.
-Results show that Judge achieves 93% consistency with expert judgments, with LLM providing logically coherent Chain-of-Thought(CoT) rationales.
-Based on these findings, we conclude that LLM-as-a-Judge represents the optimal evaluation approach for this task.
+Results show that Judge can provide logically coherent Chain-of-Thought(CoT) rationales and it achieves 93% consistency with expert judgments.
+Based on these findings, we conclude that LLM-as-a-Judge is the optimal evaluation approach for this task.
 As demonstrated in Figure 4, we reconstruct the dialogue pipeline from BERT-based DC-ICSF architecture to LLM-based Function Call architecture.
-Evaluated through LLM-as-a-Judge, the new architecture improves accuracy from 81.43% to 94.74% compared to the legacy system.
+Evaluation through LLM-as-a-Judge demonstrates that the new architecture improves accuracy from 81.43% to 94.74% compared to the original one.
 3.2 Knowledge Transfer
@@
 Building on our agent-oriented large model design, we propose Supervised Fine-Tuning (SFT) on smaller models using large model-generated data instead of relying solely on in-context learning (ICL) with large models.
-This approach offers two key benefits: (1) Smaller models can achieve performance comparable to large models when trained on sufficient high-quality data; (2) It significantly boosts inference and optimization efficiency.
-For a dataset $D$ containing prompt component $x$ and corresponding response component $y$, we define $D(x,y)=\{x_{i},y_{i}\}_{i=1}^{N}$, where $N$ denotes the dataset size.
+This approach offers two key benefits: (1) Smaller models can achieve performance comparable to large models after trained on sufficient high-quality data; (2) It significantly boosts inference and optimization efficiency.
+For a dataset $D$ containing prompt component $x$ and response component $y$, we define $D(x,y)=\{x_{i},y_{i}\}_{i=1}^{N}$, where $N$ denotes the dataset size.
 To minimize the Cross Entropy (CE) loss, we first utilize the SFT optimization, and then employ further optimization Rejection sampling Fine-Tuning (RFT).
@@
 Reinforcement Learning
-Since OpenAI introduced Reinforcement Learning from Human Feedback (RLHF) [61], it has become an essential component for aligning LLMs, as evidenced in major technical reports [62; 63; 64; 11; 65].
-DeepSeek-R1 [43] proposed a cost-effective RL training method from scratch which combined rule-based ORM and GRPO [47].
-Subsequent works like Logic-RL [8; 9; 10] also successfully applied this efficient approach across various tasks.
-We adapt this method to allow the model to automatically generate Chain of Thought (CoT) during training, which enhances performance [66] and provides insights into model reasoning for subsequent optimization.
-Also, we primarily follow Logic-RL [9], refining the reward design for this specific task (see Appendix B) while utilizing the same training dataset as RFT to achieve enhanced model performance.
+Since OpenAI introduced Reinforcement Learning from Human Feedback (RLHF) [65], it has become an essential component for aligning LLMs, as evidenced in major technical reports [66; 67; 68; 11; 69].
+DeepSeek-R1 [43] proposes a cost-effective RL training method from scratch which combined rule-based ORM and GRPO [47].
+Subsequent works like Logic-RL [8; 9; 10] applies this efficient approach across various tasks, enabling the model to automatically generate Chain of Thought (CoT) during training, thereby enhancing performance [70].
+Based on our observation that model generalization degrades after RFT, we attempt to employ RL to improve model performance on out-of-domain datasets.
+Following Logic-RL [9], we refine the reward design for our specific task (see Appendix B) and conduct extensive attempts to integrate knowledge distillation (see Appendix K).
 Distillation
@@
 Pruning
-To further compress model size, we design a two-stage pruning pipeline following NVIDIA [67], Llama3.2 [68], and Apple [69].
-Offline One-Shot Struct Pruning: Following Minitron[31], we explore two pruning strategies, for Depth Pruning, we compute layer importance scores and remove the $N$ least important layers based on the target pruning ratio, for Width Pruning, we conduct a model architecture search using PPL-based criteria and prune channels according to group-wise importance rankings, with full implementation details in Section I.1.
-Recovery training: After offline one-shot pruning, we execute a two-phase recovery process: (1) General Instruction Fine-tuning via Knowledge Distillation: Adopting Minitron’s approach, we leverage logits from a larger general instruction teacher model for recovery training.
-Due to unavailability of the teacher’s original training data, we first perform model correction on our selected IFT dataset to stabilize distillation, followed by top-k logits distillation(Section 3.2); (2) Domain-Specific Task Alignment: The pruned compact LM from phase 1 undergoes further aligning through knowledge distillation, producing a final domain-optimized LLM.
+To further compress model, we design two phases of pruning following NVIDIA [71], Llama3.2 [72], and Apple [73]: (1) Offline One-Shot Struct Pruning: Following Minitron[31], we explore two pruning strategies.
+For Depth Pruning, we compute layer importance scores and remove the $N$ least important layers based on the target pruning ratio.
+For Width Pruning, we conduct a model architecture search using PPL-based criteria and prune channels according to group-wise importance rankings, with full implementation details in Section I.1; (2) Recovery training: The first step is General Instruction Fine-tuning via Knowledge Distillation.
+Adopting Minitron’s approach, we leverage logits from a larger general instruction teacher model for recovery training.
+Due to unavailability of the teacher’s original training data, we first perform model correction on our selected IFT dataset to stabilize distillation, followed by top-k logits distillation(Section 3.2).
+The second step is Domain-Specific Task Alignment.
+The pruned compact LM undergoes further aligning through knowledge distillation, producing a final domain-optimized LLM.
 Quantization
-Additionally, quantization significantly reduces GPU memory consumption and enhances inference performance.
-In this work, we apply post-training quantization to the smaller models obtained through distillation/pruning.
+Quantization typically reduces GPU memory consumption and enhances inference performance.
+In this work, we apply post-training quantization to smaller models obtained through distillation/pruning.
 Specifically, we implement two quantization approaches: 4-bit weight-only GPTQ algorithm [33] and full 8-bit weight-activation FP8 quantization [41].
 The 4-bit weight-only quantization [33] compresses linear weights from FP16 to per-group INT4 via integer quantization, which are dequantized to FP16 during computation.
-This method achieves notable inference acceleration in memory-bound scenarios with large model parameters and low GPU memory bandwidth.
+This approach achieves notable inference acceleration in memory-bound scenarios with large model parameters and low GPU memory bandwidth.
 The FP8 quantization [41] quantizes both weights and activations to 8-bit floating-point format.
-Compared to 8-bit integer quantization [41], FP8 demonstrates superior precision and leverages NVIDIA’s FP8 tensor cores for effective inference acceleration [70].
+Compared to 8-bit integer quantization [41], FP8 shows a more noticeable improvement in accuracy, and by utilizing NVIDIA’s FP8 tensor cores, FP8 can achieve effective inference acceleration [74].
 4 Experiments
@@
 Datasets
-We conduct experiments using a large-scale dataset developed by our team, leveraging years of experience and practice in the field.
-The dataset transforms a real system’s intent recognition and semantic slot framework into an API document specification, creating a function pool with 74 functional modules and 225 parameters.
-We sample 200,000 data points from a single business scenario as the training set, with corresponding labels synthesized based on the large language model (LLM) prototype outlined in Section 3.1.
-Additionally, we employ a combination of hierarchical and random sampling to generate 11,922 samples as an in-domain test set, and 7,867 samples from a different business scenario as an out-domain test set, where some callable functions are absent from the training set.
-Compared to open-source datasets, our custom dataset is more intricate and closely mirrors real-world scenarios (see Table 1).
+In our experiments, we utilize a large-scale dataset built from years of team experience and practice accumulated in the field.
+We reconstruct the intent classification and semantic slot framework within the real-world system into an API document specification, creating a function pool comprising 74 functional modules and 225 parameters.
+We sample 200k data points from a single business scenario as the training set and synthesize labels based on the LLM prototype outlined in Section 3.1.
+A combination of stratified and random sampling yields 11,922 samples from this scenario’s data as the in-domain test set.
+Additionally, we sample 7,867 data points using the same methodology from another business scenario, without some callable functions in the training set, as the out-of-domain test set.
+Compared to open-source datasets, our self-constructed dataset is more complex and closely aligned with real-world scenarios (see Table 1).
 Metrics
@@
 Training
-We select small models (0.5B-7B) from the Qwen2.5 model family [11] for RFT, perform RL on the 3B model, and pruning on the 0.5B model, obtaining Qwen2.5-0.4B-Width and Qwen2.5-0.4B-Depth, both achieved under a 20% pruning ratio.
-Then, we perform KD on the pruned models and 0.5B model with the 7B RFT-trained teacher.
-Finally, we apply model quantization using W8A16/W4A16/W8A8 strategies to the KD-trained 0.5B model, detailed in Appendix E.
+During the training phases, base models are all from small instruct models (0.5B-7B) of the Qwen2.5 model family [11].
+We conduct RFT on them and RL on 7B models.
+We perform pruning on 0.5B models and obtain Qwen2.5-0.4B-Width and Qwen2.5-0.4B-Depth, both under a 20% pruning ratio.
+Then, we transfer knowledge from 7B RFT-trained models to pruned and 0.5B models via KD.
+We also apply KD from 7B RL-trained models to 0.5B models, followed by another RL phase (denoted as KD*+RL).
+Finally, we use model quantization via W8A16/W4A16/W8A8 strategies on the KD-trained 0.5B model, detailed in Appendix E.
 The training process is cost-efficient, utilizing only 8 NVIDIA H20 GPUs.
@@
 Evaluation settings
-We evaluate our models with the BERT-Based System, which consists of two models based on BERT-Large (330M)[71].
+We do evaluation by comparing our models with crowdsourced labelers and the BERT-based system consisting of two models based on BERT-large (330M)[75].
 To evaluate the effectiveness of our method, we invite expert annotators alongside the LLM-as-a-judge-based metric AR (see Equation 6) for performance assessment.
-We also implement a retrieval module to select the top 10 functions from the function pool, aiming to accelerate both the training and inference phases, as detailed in Appendix D.
+We also implement a retrieval module to select top 10 functions from the function pool, aiming to accelerate both training and inference phases, as detailed in Appendix D.
 In real-world scenarios, we deploy models on NVIDIA L20 GPU and perform effectiveness evaluation via TensorRT-LLM[74], with specific configuration details provided in Appendix H.
-During the comparison between the trained models and a BERT-based system, we adopt queries per second (QPS) as the primary metric for system throughput, response time (RT) for end-to-end latency, and concurrent clients (CC) to quantify high-concurrency performance.
+During the comparison between the trained models and the BERT-based system, we adopt queries per second (QPS) as the primary metric for system throughput, response time (RT) for end-to-end latency, and concurrent clients (CC) to quantify high-concurrency performance.
 4.2 Experimental Results
-Table 2: Main results of the models trained on Qwen2.5.(See Appendix J for FLOPs analysis)
+Table 2: Main results of the models trained on Qwen2.5(See Appendix J for FLOPs analysis).
 Model / System AR GFLOPs
 Crowdsourced Labelers 77.85% -
-BERT-Based System 81.43% 38.79
-Qwen2.5-72B (ICL) 94.74% 5223.91
-Qwen2.5-7B (RFT) 95.29% 572.47
-Qwen2.5-3B (RFT) 95.30% 297.98
-Qwen2.5-1.5B (RFT) 95.16% 171.19
-Qwen2.5-0.5B (RFT) 94.66% 77.89
-Qwen2.5-0.5B (KD) 94.85% 77.89
+BERT-based System 81.43% 38.65
+Qwen2.5-72B (ICL) 94.74% 4907.50
+Qwen2.5-7B (RFT) 95.29% 434.04
+Qwen2.5-3B (RFT) 95.30% 218.94
+Qwen2.5-1.5B (RFT) 95.16% 111.91
+Qwen2.5-0.5B (RFT) 94.66% 43.31
+Qwen2.5-0.5B (KD) 94.85% 43.31
 Qwen2.5-0.5B-W8A16 (KD) 94.76% -
@@
 Qwen2.5-0.5B-W8A8 (KD) 94.83% -
-Qwen2.5-0.4B-Width (KD) 94.17% 72.93
-Qwen2.5-0.4B-Depth (KD) 94.20% 67.68
+Qwen2.5-0.4B-Width (KD) 94.17% 33.10
+Qwen2.5-0.4B-Depth (KD) 94.20% 40.83
 Figure 3: Cost and performance results. FLOPs reduced by 10x with nearly lossless AR.
 Table 3: RFT vs RL on in-domain data and out-domain data.
-Model In-
-Domain
-Out-
-Domain
+Model AR
 AR
 Qwen2.5-72B (ICL) 94.74% 94.63%
-Qwen2.5-0.5B (RFT) 94.66% 89.32%
-Qwen2.5-3B (RFT) 95.30% 91.02%
-Qwen2.5-3B (RL) 94.68% 92.95%
+Qwen2.5-7B (RFT) 95.29% 90.86%
+Qwen2.5-0.5B (RFT) 94.66% 90.62%
+Qwen2.5-7B (RL) 95.36% 94.80%
+Qwen2.5-0.5B (KD*+RL) 95.55% 91.68%
 Figure 4: Comparison over different models by professional annotators
 LLM Prototype Preferred by Both Human and LLM-as-a-Judge
-As illustrated in Figure 4, annotations provided by professional annotators reveal that the LLM prototype (72B ICL) achieves a win+tie rate of 95.59% over the BERT-based system and 92.98% over the crowdsource labeler.
+As illustrated in Figure 4, annotations provided by professional annotators reveal that the LLM prototype (72B ICL) achieves a win+tie rate of 95.59% over the BERT-based system and 92.98% over crowdsourced labelers.
 This demonstrates the superior performance of the LLM prototype.
-Using the LLM-as-a-Judge for evaluation, we find that the LLM Prototype surpasses the Bert-Based System by 13.31% and crowdsourced annotations by 16.89%, as displayed in Table 2.
-This conclusion is roughly consistent with those provided by professional annotators, demonstrating the effectiveness of the LLM judge.
+Through the evaluation via LLM-as-a-Judge, we find that the LLM Prototype surpasses the BERT-based System by 13.31% and crowdsourced annotations by 16.89%, as shown in Table 2.
+This conclusion is nearly consistent with that presented by professional annotators, demonstrating the effectiveness of the LLM judge.
 RFT and KD as Effective Knowledge Transfer Methods
 As demonstrated in Table 2, models trained with RFT generally exhibit a higher AR than the LLM prototype.
-Specifically, both the 3B and 7B models display negligible differences in performance but each achieve a 0.5% improvement in AR compared to the LLM prototype.
-In contrast, the 0.5B model shows a 0.5% lower AR than the 1.5B model.
-Moreover, KD emerges as a valuable complement to RFT, enhancing the performance of the 0.5B model by an additional 0.19%.
+Specifically, both the 3B and 7B models display negligible differences in performance and both achieves a 0.5% improvement in AR compared to the LLM prototype.
+The 0.5B model’s AR is 0.5% lower than the 1.5B model, while KD emerges as a valuable complement to RFT, enhancing the performance of the 0.5B model by an additional 0.19%.
 In terms of inference performance, the 0.5B model demonstrates only a slight 0.4% difference in AR compared to the 7B model, while achieving a significant 13.57x improvement in QPS and a substantial 91.3% reduction in latency, as illustrated in Table 4.
 RL’s Superior Generalization Ability
-According to Table 3, while RFT performs well with in-domain data, it demonstrates poorer results with out-of-domain data, showing a 3.61% AR decrease compared to the LLM prototype.
-Conversely, RL models present a smaller gap (1.68%) with the LLM prototype on out-of-domain data, indicating better generalization of RL training methodologies.
-Tiny LMs as a Cost-Efficient Upgrade to BERT-based System
-Table 2 and Table 4 illustrate that quantization introduces a negligible AR loss of less than 0.1% while significantly enhancing inference performance.
-In the optimal scenario, quantization results in a 1.29x increase in QPS and a 0.76x decrease in RT.
-On the other hand, pruning results in a slightly higher accuracy loss, approximately 0.2%, but offers a more consistent 20% improvement in performance.
-In terms of end-to-end inference performance, under the condition where the CC is 64, the BERT-based system achieves a QPS of 1110.88 and an RT of 55.84 ms.
-Despite the cost being approximately doubled compared to the BERT-based system, the AR is improved by an absolute increase of 13%.
+According to Table 3, while RFT performs well on in-domain data, it demonstrates worse results on out-of-domain data with a 3.61% AR decrease compared to the LLM prototype.
+Conversely, 7B RL-trained models achieve superior performance compared to other models on both in-domain and out-of-domain datasets, indicating better generalization of RL training methodology.
+Notably, even with models as small as 0.5B, employing a hybrid training methodology with KD+RL can yield superior in-domain results and an improvement of 1% in out-of-domain datasets compared to the RFT method.
+For more detailed exploration of the RL results, please refer to Appendix K.
+Tiny LLMs as a Cost-Efficient Upgrade to BERT-based System
+As shown in Table 2 and Table 4, quantization minimally impacts accuracy, with less than a 0.1% AR loss, while significantly enhancing inference performance, achieving up to 1.29x higher QPS and 0.76x lower RT in the optimal scenario.
+Pruning incurs a slightly higher 0.2% accuracy loss but consistently enhances performance by 20%.
+For end-to-end inference, with a CC of 64, the BERT-based system reaches 1110.88 QPS and 55.84 ms RT.
+Although the cost is roughly twice that of the BERT-based system, the AR improves by up to absolute 14%.
 Table 4: Inference Performance under different model compression techniques.
@@
 Our experimental analysis highlights a compelling trade-off between model size and computational efficiency.
-Although larger models typically demonstrate superior performance, the expected disparity in performance between large and smaller models is not as pronounced as anticipated.
-As shown in Figure 3, the 0.4B model achieves performance almost on par with the 7B model, efficiently requiring 11.8% FLOPs, while even reaches more than 17.4x QPS improvement and 92.9% RT reduction, with the 7B model’s RT being 882.552 ms under 16 CC.
-Although this model has a slight gap in cost compared to traditional BERT models, the attractive accuracy and scalability stemming from LLM base makes it a favorable choice for various applications.
-Characteristics of Training Paradigm
+Although larger models typically show superior performance, the expected disparity between large and small ones is not as pronounced as anticipated.
+As shown in Figure 3, the 0.4B model achieves performance almost on par with the 7B model, efficiently requiring 11.8% FLOPs and reaching more than 17.4x QPS improvement and 92.9% RT reduction, with the 7B model’s RT being 882.552 ms under 16 CC.
+Although this model has a slight gap in cost compared to traditional BERT models, the improved accuracy and scalability stemming from LLM base makes it a favorable choice for various applications.
+Characteristics of Training Paradigms
 The results presented in Table 3 underscore the distinct characteristics of different training methodologies.
-Fine-tuning primarily focuses on memorizing data, making it susceptible to overfitting to the training set distribution.
-In contrast, RL demonstrates greater robustness by not merely fitting individual data points.
+Fine-tuning primarily focuses on memorizing data, making it susceptible to overfitting to the training set.
+In contrast, RL presents greater robustness by not merely fitting individual data points.
 We posit that RL’s superior generalization capabilities stem from its ability to learn from negative samples, which are generated by the model being trained itself.
-RL is highly effective for dealing with well-structured problems, while it presents unsatisfactory convergence facing ambiguous data or a lack of domain knowledge in real-world tasks.
-Moreover, RL is sensitive to noisy data.
-Despite these limitations, RL remains a promising direction for development, especially when building upon the foundation of traditional fine-tuning training strategies.
+RL is highly effective for dealing with well-structured problems, while it has unsatisfactory convergence facing ambiguous data or a lack of domain knowledge in real-world tasks.
+Additionally, applying RL directly to tiny models often proves to be ineffective, likely due to such models’ limited capacity to utilize Chain-of-Thought (CoT) rationales.
+In such cases, KD can compensate for this limitation by training the tiny models on a large amount of CoT data generated by a RL-trained teacher model.
+Compared to fine-tuning, KD can better extract CoT rationales from the probability distribution.
+Consequently, the hybrid methodology of KD and RL can achieve remarkably enhanced performance and represents a promising direction for future development.
 Comparison over Model Compression Strategies
 Our analysis identifies depth pruning as the optimal approach for domain-specific models.
-Depth pruning achieves 20% inference acceleration with lossless AR at equivalent compression ratios (Table 2), stemming from the nature of depth pruning, removing entire transformer layers and their associated attention computations.
-Although width pruning exhibits slightly better effects in general instruction tasks after training (Table 9), depth pruning demonstrates optimal efficiency facing long prompts, with width pruning focusing on matrix optimizations (Appendix J).
-Regarding quantization, the W4A16 strategy achieves impressive performance in single-thread scenarios, while FP8 outperforms under high concurrency conditions with lossless accuracy, considered a preferred deployment paradigm, highlighting the importance of throughput.
+Depth pruning achieves 20% inference acceleration with lossless AR at same compression ratios (Table 2), stemming from the nature of depth pruning, removing entire transformer layers and their associated attention computations.
+Although width pruning exhibits slightly better effects in general instruction tasks after training (Table 9), depth pruning presents optimal efficiency facing long prompts, with width pruning focusing on matrix optimizations (Appendix J).
+Regarding quantization, the W4A16 strategy achieves impressive performance in single-thread scenarios, while FP8 outperforms with lossless accuracy under high concurrency conditions as the superior paradigm for throughput deployments.
 6 Conclusion
@@
 It advocates for a balanced research agenda that equally emphasizes foundational advancements and domain-specific applications.
-Cost-Efficient Three-Stage Pipeline
-The proposed pipeline—comprising prototyping, knowledge transfer, and model compression—presents a systematic framework designed to achieve high performance while maintaining low costs in both training and inference.
+Scalable Cost-Efficient Three-Stage Pipeline
+The proposed pipeline-comprising prototyping, knowledge transfer, and model compression-presents a systematic framework designed to achieve high performance while maintaining low costs in both training and inference.
 We believe that this approach can be effectively adapted to other domains, thereby potentially advancing research in those areas and facilitating the widespread deployment of LLMs.
 7 Limitation & Future Works
-While our proposed three-stage pipeline demonstrates promising results in cost-performance, several limitations warrant further investigation, underscoring the need for holistic optimization across algorithmic robustness, dynamic adaptation and hardware efficiency to enable truly scalable and trustworthy adoption in industrial settings:
-Continual Learning for Dynamic Adaptation
-While cost-efficient, the static training-deployment cycle of our pipeline hinders its adaptation to dynamic tasks and evolving datasets.
-Future work should develop advanced continual learning frameworks that enable incremental knowledge integration without catastrophic forgetting, leveraging rapid training cycles to maintain deployment agility.
-Employing online learning and lifelong learning will enhance models’ robustness and adaptability, ensuring sustained relevance and precision in changing linguistic contexts and shifting user demands.
-Edge Deployment Optimization with Performance Assurance
-Being sufficiently compact, our super tiny models have demonstrated impressive capabilities in currently high-demand applications of IoT devices, such as mobile phones and tablets, characterized by low latency, cost-effectiveness, cloud independence, and robust privacy.
-While further compression to an even smaller size like 0.1B may not be significantly beneficial, exploration of optimizing on-device deployment of such ultra-small models remains a promising direction for future research.
-Future works should develop hardware-aware compression algorithms and adaptive runtime kernels to optimize latency-energy tradeoffs across diverse IoT ecosystems, addressing real-world heterogeneous hardware constraints such as memory fragmentation and NPU compatibility.
-We will also explore more resource-efficient deployment techniques for edge devices—such as sparse computation, edge-centric quantization, and federated learning.
+Our pipeline has presented optimal cost-performance results, and our super tiny models have exhibited remarkable capabilities, featuring low latency, cost efficiency, cloud independence, and robust privacy.
+Further compression may offer limited marginal benefits.
+Thus, we are more focused on enhancing such models’ capabilities and expand their broader applications such as mobile devices:
+Hybrid Methodology for LLMs Capabilities Transfer
+Based on our existing experiments, the mixer of reinforcement learning and knowledge distillation has effectively enhanced the generalization capabilities of super tiny models.
+This implies a potentially powerful hybrid methodology that allows super tiny models to nearly fully retain the capabilities of LLMs in domain-specific tasks.
+In future research, we will build upon our previous work to conduct deeper studies and theoretical analyses of this hybrid approach in transferring LLMs capabilities to such super tiny models.
+On-device Deployment Optimization
+Although we achieve 8-bit compression and online deployment, real-world edge devices face heterogeneous hardware constraints such as memory fragmentation and NPU compatibility.
+In the future, for on-device deployment of super tiny models, we will develop hardware-aware compression algorithms and adaptive runtime kernels to optimize latency-energy tradeoffs across diverse IoT ecosystems, and explore more resource-efficient deployment techniques for edge devices, such as sparse computation, edge-centric quantization, and federated learning.
 Appendix A Dataset Analysis
-Type-Token Rate (TTR) is a fundamental measure of lexical diversity that reflects the variability of vocabulary locally, but it is sensitive to text length. The calculation formula is:
+Type-Token Rate (TTR), as a fundamental measure of lexical diversity, reflects the local variability of vocabulary, but it is sensitive to text length. The calculation formula is:
 $TTR=\frac{N_{types}}{N_{tokens}}$ (7)
 where $N_{tokens}$ denotes the total number of 2-gram patterns in the text, and $N_{types}$ represents the number of distinct 2-gram patterns.
-Entropy is an information-theoretic metric that quantifies the uncertainty within binary grammatical sequences, based on their probability distribution.
+Entropy, as an information-theoretic metric, quantifies the uncertainty within binary grammatical sequences based on their probability distribution.
 A higher entropy value indicates more complex local language patterns.
@@
 where $p_{i}$ is the probability of occurrence for the $i^{th}$ 2-gram pattern.
-Both TTR and Entropy are sensitive to text length.
-To enable a relatively fair comparison of the complexity across datasets, we randomly sample 700 entries from the test set of each dataset for the computation(as the ATIS dataset only contains 700 instances).
+Since both TTR and Entropy are sensitive to text length, to ensure a relatively fair comparison of dataset complexity, we randomly sample 700 entries from the test set of each dataset for calculation(as the ATIS dataset only contains 700 instances).
 Appendix B Details for Reinforcement Learning
 Our reward design consists of two components, Format Reward($S_{format}$) and Answer Reward ($S_{answer}$):
-  • For Format Reward, we adopt the logic of Logic-RL[9], which utilizes regular expressions to validate the output format against the specified template <think>...</think><answer>...</answer>.
-Additionally, we extend this design by verifying the legality of function usage and its parameters.
-For example, if it choose a function that doesn’t exist in the input function pool, then it is considered as invalid.
+  • For Format Reward, we adopt the logic of Logic-RL[9], which validates the output format against the specified template <think>...</think><answer>...</answer> via regular expressions.
+We add additional designs to ensure the validity of the usage of the function and its parameters.
+If a function not present in the predefined input function pool is selected, it is deemed invalid.
 $S_{format}=\begin{cases}1,&\text{if the format is correct, and the function call is valid}\\ 0,&\text{otherwise}\end{cases}$ (9)
-  • For Answer Reward, according to our specific scenario, we find that the model attempts to directly guess answer during the training, leading no association between the Chain-of-Thought (CoT) and the final results.
+  • For Answer Reward, in our specific scenario, we observe that the model tends to directly guess answer during training, resulting in no association between the Chain-of-Thought (CoT) and the final results.
 Requiring explicit repetition of the answer within the CoT framework effectively resolves this issue.
-As demonstrated in Figure 6, we optimize its original design in order to make the Chain-of-Thought (CoT) reasoning and the results more consistent.
-We require: (1) the CoT must explicitly reference the target function at least once; (2) the final function mentioned in the CoT must match the function in the answer.
-Otherwise, it will be considered as a parsing error.
-$S_{answer}=\begin{cases}2,&\text{if the answer is correct,}\\ &\text{and the thought ends with the function in the answer,}\\ -2&\text{if the answer cannot be parsed or is missing,}\\ &\text{or is not defined in the function list,}\\ &\text{or the thought doesn't end with the function in the answer}\\ -1.5,&\text{otherwise}\end{cases}$ (10)
+As illustrated in Figure 6, we optimize its original design in order to enhance the consistency between the Chain-of-Thought (CoT) reasoning and the results.
+Our requirements are as follows: (1) The CoT must explicitly reference the target function at least once; (2) The final function mentioned in the CoT must correspond to the function in the answer.
+Failure to comply with these conditions will be regarded as a parsing error.
+Additionally, we incorporate fine-grained rewards, which provide better guidance when the answer is partially correct, enabling the model to achieve improved performance.
+$S_{answer}=\begin{cases}-2&\text{if the answer is missing, or is unable to be parsed,}\\ &\text{or is not defined in the function list,}\\ &\text{or the thought doesn't end with the function in the answer}\\ -1.25&\text{if the names of at least one parameter and function are correct,}\\ -1.2&\text{if the name of function, as well as the name }\\ &\text{and value of at least one parameter are correct}\\ 2,&\text{if the answer is correct,}\\ -1.5,&\text{otherwise}\end{cases}$ (10)
 Figure 6: An example of optimization of the original Answer Reward design by requiring explicit repetition of the answer within the CoT framework.
@@
 Appendix D Retriever Setup
-Initially, we employ GTE-large [59] as the embedding model trained based on the BERT-large [71].
-We compare the cosine similarity of the embedding between the API description generated by LLM based on this model and the API description in the API document, but the recall accuracy is not high.
-Subsequently, we replace it with GTE-Qwen2-1.5B-instruct [59] trained based on Qwen2-1.5B [97]. This model has the advantage of being able to fine tune matching tasks by modifying prompts.
-Next, based on this model and a customized prompt in Section D.1, we achieve a higher accuracy by matching the cosine similarity between the API description generated by the model and the JSON text of the entire API document.
-During training, we optimize the API description model through rejection sampling, with implementation details provided in Section D.2.
-Finally, to achieve maximal inference speed, we trained a classifier based on BERT-large [71] for top-1 recall using rejection sampling.
-The classifier takes Dialog, User Query, and necessary context as input and outputs the function category.
-Then, the functions corresponding to the top-k output logits are considered as candidate functions for the subsequent model’s function call.
-This approach significantly improves inference speed (due to the inherent efficiency of BERT classifiers compared to autoregressive GPT models) and also enhances function recall.
-However, this comes at the cost of scalability, as retraining the recall model is required for each update to the function pool.
-As illustrated in Table 5, these optimizations substantially improve API recall rate.
+Initially, we select GTE-large [63], which is trained on the BERT-large [75], as our embedding model.
+Using this model, we generate embeddings of API descriptions produced by the LLMs.
+Then, we compare these embeddings with those of the API description from the document using cosine similarity.
+However, this model’s recall accuracy is not high.
+Subsequently, we switch to GTE-Qwen2-1.5B-instruct [63], a model trained on Qwen2-1.5B [76], which offers the advantage of fine-tuning for matching tasks through prompt modification.
+Next, utilizing this model with a customized prompt in Section D.1, we perform cosine similarity matching between the model-generated API descriptions and the JSON text of the entire API documentation, achieving a higher accuracy.
+During training, we optimize the API description model using rejection sampling, with implementation details provided in Section D.2.
+Finally, to achieve optimal inference acceleration, we train a classifier based on BERT-large [75] for top-1 recall using rejection sampling.
+The classifier takes Dialog, User Query, and necessary context as input and the function category as output.
+The functions corresponding to the top-k output logits are then considered as candidate functions for the subsequent model’s function call.
+This approach significantly accelerates inference due to the efficiency of BERT classifiers compared to autoregressive GPT models, while also enhancing function recall capabilities.
+However, it introduces scalability challenges since the recall model needs to be retrained with each update to the function pool.
+As illustrated in Table 5, this series of optimizations lead to a substantially improvement in API recall rates.
 Table 5: Comparison of retrieval performance of models: GTE-large model, GTE-Qwen2-1.5B-instruct model, optimized API description model + GTE-Qwen2-1.5B-instruct model.
@@
 {mdframed}
-Instruction: Given a API description, retrieve the most similar API from the function pool
+Instruction: Given a API description, retrieve the most similar API from the function pool.
 Query:
@@
 Appendix E Training Details
-Table 6: Training Parameters in different stages.
+Table 6: Training Parameters in different phases.
 Supervised Fine-tuning (SFT)
@@
 […]
-The main training parameters are shown in Table 7.
-During the training phase, we utilize the AdamW optimizer [96], setting the model’s maximum sequence length to $5120$.
-In the RL phase, we employ the baseline variant of Reinforce++ [50] as the reinforcement learning algorithm, offering advantages of efficiency, conciseness, and training stability.
+The main training parameters are detailed in Table 7.
+During the training phase, we employ the AdamW optimizer [77], setting the model’s maximum sequence length to $5120$.
+In the RL phase, we utilize the baseline variant of Reinforce++ [50] as the reinforcement learning algorithm, offering advantages in efficiency, conciseness, and training stability.
 Following [78], the K2 estimator is used instead of the original K1 in Reinforce++.
 In the AKL phase, the threshold parameter $\mu$ is reduced from 1 to 0.5.
-For pruning, we use the 0.5B-Instruct model as the base model and employ width/depth pruning strategies with 20%/30% prune ratio, finally obtaining two models Qwen2.5-0.4B(0.35B)-Width and Qwen2.5-0.4B(0.35B)-Depth.
-We produce the extremely compressed models through a two-stage training process (accuracy recovery training, KD).
-Complete training details for the pruned models can be found in Appendix I.
-For quantization, we primarily experiment with three quantization strategies on the 0.5B model trained with AKL: W8A16, W4A16, and W8A8.
-Both W8A16 and W4A16 adopt per-group symmetric int8 weight-only quantization with a group_size=128, and activations are in fp16.
+For pruning, we use the 0.5B-Instruct model as the base, implementing width/depth pruning strategies with a 20%/30% prune ratio, resulting in two models Qwen2.5-0.4B(0.35B)-Width and Qwen2.5-0.4B(0.35B)-Depth.
+These extremely compressed models are produced through a two-phase training process involving accuracy recovery training and knowledge distillation (KD).
+Comprehensive training details for the pruned models are available in Appendix I.
+For quantization, we primarily experiment with three strategies on the 0.5B model trained with AKL: W8A16, W4A16, and W8A8.
+Both W8A16 and W4A16 employ per-group symmetric int8 weight-only quantization with a group_size=128, while activations are in fp16.
 To improve accuracy, W4A16 use 512 training data samples combined with the GPTQ algorithm [33] for calibration.
-In contrast, W8A8 applies FP8 for both weights and activations.
-For the training frameworks, we use Megatron-LM [73] for SFT and KD, while OpenRLHF [74] is employed for RL.
-Data sampling during training is implemented using vLLM [72].
-In the evaluation phase, we utilize SGLang [75] to generate data.
+In contrast, W8A8 uses FP8 for both weights and activations.
+For the training frameworks, we utilize Megatron-LM [79] for SFT and KD, while OpenRLHF [80] is employed for RL.
+Data sampling during training is conducted using vLLM [81].
+In the evaluation phase, we employ SGLang [82] to generate data.
 Appendix F Computing resources
-We utilize 8 NVIDIA H20 GPUs for training process and data generation.
-For inference performance evaluation, we employ a single NVIDIA L20 GPU.
-We apply an INTEL XEON PLATINUM 8575C processor with 192 cores for computation, which operates at a frequency of 3028.956 MHz and a cache size of 327680 KB.
-In addition, we establish a virtual development environment based on Kubernetes, leveraging 64 virtual cores and 512 GB of memory.
-During the training phase, the RFT process takes approximately 2 to 14 hours for models ranging from 0.5B to 7B.
-Pre-KD logits caching for a 7B model requires about 5 hours, while RL for the 3B model takes approximately 3 days.
-The KD process for the 0.5B model is completed in about 2 hours.
-Training each pruned model demands roughly 3.5 days.
-Furthermore, for the LLM prototype and LLM-as-a-Judge, data generation of every 10,000 entries requires approximately 1 hour.
+We utilize 8 NVIDIA H20 GPUs for training and data generation processes.
+For evaluating inference performance, a single NVIDIA L20 GPU is employed.
+We apply an INTEL XEON PLATINUM 8575C processor with 192 cores for computation, operating at a frequency of 3028.956 MHz and a cache size of 327680 KB.
+Additionally, we establish a virtual development environment based on Kubernetes, leveraging 64 virtual cores and 512 GB of memory.
+During the training phase, the RFT process spans approximately 2 to 14 hours for models ranging from 0.5B to 7B in size.
+Pre-KD logits caching for a 7B model requires about 5 hours, whereas RL for the 3B model extends to approximately 3 days.
+The KD process for the 0.5B model is completed within about 2 hours, and training each pruned model requires roughly 3.5 days.
+Moreover, for the LLM prototype and LLM-as-a-Judge, generating data for every 10,000 entries takes approximately 1 hour.
 Appendix G Detailed Analysis for Experimental Results for KD
-As demonstrated in Table 7, it is evident that balancing FKL and RKL is quite crucial for our task.
-The process of function selection by the model resembles a classification task, whereas the generation of parameter names and values for function calls is more akin to a generation task.
+As demonstrated in Table 7, it is clear that balancing FKL and RKL is crucial for our task.
+The function selection process by the model resembles a classification task, whereas generating parameter names and values for function calls is more akin to a generation task.
 The parameter $\mu$ in AKL is precisely used to adjust the effects between classification and generation.
 Experimental results demonstrate that at $\mu\leq 0.7$ and $\mu=1.0$ (equivalent to FKL), the performance with AKL is similar to or worse than RFT.
-Based on our findings where the average probability of the top-1 token from the teacher model exceeds 99%, we draw the conclusion that setting a lower $\mu$ value causes the effect of AKL to excessively shift more towards RKL.
-As described in the AKL paper, RKL prioritizes learning the tail of the probability distribution, which is detrimental for rapid convergence in scenarios without requiring diversity.
+Based on our findings where the average probability of the top-1 token from the teacher model exceeds 99%, we conclude that setting a lower $\mu$ value causes the effect of AKL to excessively shift more towards RKL.
+As described in the AKL paper, RKL prioritizes learning the tail of the probability distribution, which can impede rapid convergence in scenarios without requiring diversity.
 However, RKL can bring considerable benefits for the function classification part.
-Overall, AKL demonstrates greater potential compared to RFT, achieving some performance improvements with an appropriately chosen $\mu$ value.
+Overall, AKL demonstrates greater potential compared to RFT, achieving performance improvements with an appropriately chosen $\mu$ value.
 Appendix H Inference Details
 The inference framework utilizes TensorRT-LLM[74] v0.16.0 with enabled optimizations including paged attention, prefix cache, CUDA graph size set to 1024, and max token number set to 32768 for inference acceleration.
-Appendix I Prune Detail
+Appendix I Pruning Details
 I.1 Offline One-shot Pruning
-Importance Metrics for Structural Pruning
-(1) Depth Pruning: we propose Layer Importance (LI) metric based on the similarity between consecutive layers’ activation. Formally, we define the importance of the $i$-th layer as:
-$LI_{i}=1-E_{X,t}\frac{X_{i,t}^{T}X_{i+1,t}}{\|X_{i,t}\|_{2}\|X_{i+1,t}\|_{2}}$ (11)
-where $X_{i}$ denotes the activation matrix of layer $i$, with $X_{i,t}$ representing the $t$-th column (embedding vector) of $X_{i}$.
-After we collecte all layer activations $\{X_{i}\}_{i=1}^{L}$ through a single forward pass on calibration data, then remove the $N$ least important layers, where $N$ is determined by the target pruning ratio; (2) Width Pruning: We categorize width pruning along three dimensions by channel dependencies: Embedding dimension, Head dimension and Intermediate dimension(as as illustrated in ).
-For compact LLMs (e.g., Qwen2.5-0.5B with head size 14), we exclude head dimension pruning due to its demonstrated sensitivity[37], we use an activation-based importance metric Channel Importance(CI) to do Channel selection as:
+Importance as Metric for Structured Pruning
+For Depth Pruning, we adopt Layer Importance (LI) as the metric, which is based on the similarity between consecutive layers’ activation. Formally, we define the importance of the $i$-th layer as:
+$LI_{i}=1-\mathbb{E}_{X,t}\frac{X_{i,t}^{T}X_{i+1,t}}{\|X_{i,t}\|_{2}\|X_{i+1,t}\|_{2}}$ (11)
+where $X_{i}$ denotes the activation matrix of the $i$-th layer and $X_{i,t}$ denotes the $t$-th column (embedding vector) of $X_{i}$.
+After all layer activations $\{X_{i}\}_{i=1}^{L}$ through a single forward pass on calibration data are collected, the $N$ least important layers are removed, where $N$ is determined by the target pruning ratio.
+For Width Pruning, we classify width pruning into three dimensions based on channel dependencies: Embedding dimension, Head dimension and Intermediate dimension (as illustrated in Figure 8).
+For compact LLMs (e.g., Qwen2.5-0.5B with head size of 14), we exclude head dimension pruning due to its demonstrated sensitivity[37].
+Instead, we employ an activation-based importance metric, Channel Importance(CI), for channel selection as follows:
 $CI^{i}_{\text{embedding}}=\sqrt{\sum_{b,s}|\text{LN}(X)^{i}|^{2}},\quad CI^{i}_{\text{inter}}=\sqrt{\sum_{b,s}|x(W_{ffn1}^{i})^{T}|^{2}}$ (12)
-where $CI^{i}_{\text{embedding}}$ denotes the importance score of the $i$-th embedding dimension, computed from LayerNorm outputs of attention and FFN blocks, while $CI^{i}_{\text{inter}}$ represents the importance of the $i$-th intermediate dimension, calculated from the $i$-th row of FFN’s gate and up projection matrices $W_{ffn1}^{i}$.
+where $CI^{i}_{\text{embedding}}$ denotes the importance score of the $i$-th embedding dimension, computed from the LayerNorm outputs of attention and FFN blocks, while $CI^{i}_{\text{inter}}$ denotes the importance of the $i$-th intermediate dimension, calculated from the $i$-th row of FFN’s gate and up projection matrices $W_{ffn1}^{i}$.
 The importance scores are first aggregated across batch and sequence dimensions through summation $\sum_{b,s}$, then normalized using L2 normalization within each subgroup to obtain global channel importance.
-Figure 8: Pruning dependency in Qwen
-Best Width Pruning Model Configuration
-To maintain hardware-efficient matrix multiplication during inference, we align pruning granularity to 64 for Embedding dimension and 256 for intermediate dimension, yielding multiple width-pruned model configurations as shown in Table 8.
-We compute channel importance scores $CI^{(i)}_{\text{embedding}}$ and $CI^{(j)}_{\text{inter}}$ on a calibration dataset comprising 1024 randomly sampled sequences from WikiText-2.
+Figure 8: Pruning Dependency in the Qwen2.5
+Optimal Width Pruning Configuration
+To ensure hardware-efficient matrix multiplication during inference, we set the pruning granularity to 64 for embedding dimension and 256 for the intermediate dimension.
+This approach results in multiple width-pruned model configurations, as show in Table 8.
+We calculate channel importance scores $CI^{(i)}_{\text{embedding}}$ and $CI^{(j)}_{\text{inter}}$ using a calibration dataset consisting of 1024 randomly sampled sequences from WikiText-2.
 For each configuration, we prune entire channel groups with the lowest CI scores and evaluate the resulting one-shot pruned models through WikiText-2 perplexity.
-Table 8 presents the optimal configurations for width pruning at rates of 20% and 30%, which are identified as Qwen2.5-0.4B-Width-2 and Qwen2.5-0.35B-Width-3, respectively.
+Table 8 presents the optimal configurations for width pruning at rates of 20% and 30%, identified as Qwen2.5-0.4B-Width-2 and Qwen2.5-0.35B-Width-3, respectively.
 Table 8: Perplexity (PPL) on WikiText-2 for different model architectures with 20% and 30% width pruning.
+Model
 Hidden
@@
 I.2 Recovery Training in Pruning
-General Instruction Fine-tuning Dataset And Traing Hyperparameters
-We employ the Infinity-Instruct-7M[106] dataset and adopt AKL distillation[29] described in Section 3.2 for general instruction recovery training.
-We utilize Qwen2.5-7B-Instruct[11] as the teacher model, following Minitron’s methodology[67] to perform teacher correction on the Infinity-Instruct-7M dataset.
-We employ the Adam optimizer (weight decay=1e-4) with a cosine learning rate schedule decaying from 2e-5 to 2e-6.
-We conduct training for 2 epochs, utilizing a batch size of 32, and each batch is packed to accommodate sequences of up to 7168 tokens.
-Table 9: Performance of different pruned model compared to the unpruned model.
+Recovery Training Dataset and Hyperparameters
+We use the Infinity-Instruct-7M[83] dataset and apply AKL distillation[29] described in Section 3.2 for general instruction recovery training.
+Qwen2.5-7B-Instruct[11] serves as the teacher model, and we follow Minitron’s methodology[71] to perform teacher correction on the Infinity-Instruct-7M dataset.
+The Adam optimizer is used with a weight decay of 1e-4 in conjunction with a cosine learning rate schedule, decaying from 2e-5 to 2e-6.
+We conduct training over 2 epochs, utilizing a batch size of 32, with each batch packed to accommodate sequences of up to 7168 tokens.
+Table 9: Performance comparison of Qwen2.5 models after pruning.
 We report 5-shot performance on mmlu and cmmlu, 4-shot for gsm8k, zero-shot for ARC-Challenge, HellaSwag.
-* indicates that the baseline Qwen-2.5-0.5B utilizes the same instruct fine-tuning dataset for correction.
+* denotes that the baseline Qwen-2.5-0.5B utilizes the same instruct fine-tuning dataset for correction.
 Model
@@
 Qwen-2.5-0.35B-Depth 25.46 24.42 26.12 25.21 5.61 21.36
-Evaluation on Knowledge & Reasoning Tasks
-Table 9 presents the performance of pruned models on downstream tasks, including MMLU and CMMLU for knowledge assessment, as well as ARC-Challenge, HellaSwag and Winogrande for commonsense reasoning evaluation.
-Our experiments demonstrate distinct recovery patterns between width and depth pruning on general tasks.
-At 20% pruning rate, width pruning achieves 90.5% performance retention compared to the unpruned baseline, while depth pruning retains only 78.4%.
-When increasing pruning intensity to 30%, width pruning maintains 83.8% of unpruned performance, whereas depth pruning suffers catastrophic performance degradation-notably on GSM8K, where accuracy drops from 45.11 to merely 5.61.
+Performance across Knowledge Assessment & Reasoning Tasks
+Table 9 summarizes the performance of pruned models across various downstream tasks, including MMLU and CMMLU for knowledge assessment, alongside ARC-Challenge, HellaSwag and Winogrande for commonsense reasoning evaluation.
+Our experiments reveal distinct recovery patterns between width and depth pruning across general tasks.
+At a 20% pruning rate, width pruning achieves 90.5% of the performance relative to the unpruned baseline, while depth pruning retains only 78.4%.
+When increasing the pruning intensity to 30%, width pruning preserves 83.8% of the original performance.
+In contrast, depth pruning suffers catastrophic performance degradation, particularly on the GSM8K task, where accuracy drops from 45.11 to a merely 5.61.
 Ablation of Pruning Ratio on Domain-Specific Task
 Our investigation reveals a nonlinear relationship between pruning intensity and model accuracy in domain-specific tasks.
-Table 10 demonstrates that both depth-wise and width-wise pruning with 20% pruning rate incur negligible AR degradation (<0.5%).
-However, when increasing the pruning rate to 30%, we observe significant performance deterioration (>1.5% AR drop).
-These findings suggest that for small-scale models operating on domain-specific tasks, pruning rates exceeding 20% may lead to substantial accuracy compromises.
-Table 10: Results of different pruning ratios and pruning strategies on domain specific task.
+As shown in Table 10, both depth-wise and width-wise pruning at a 20% pruning rate result in a negligible <0.5% AR reduction.
+However, increasing the pruning rate to 30% leads to notable performance deterioration (>1.5% AR drop).
+These findings suggest that for small-scale models on domain-specific tasks, pruning rates exceeding 20% may significantly compromise accuracy.
+Table 10: Results with different pruning strategies on domain specific tasks.
 Model AR
@@
 For the calculation of FLOPs, we first define the notations symbolically:
-  • $B$: batch size
-  • $H$: hidden size
-  • $S$: sequence length (average number of prompt tokens)
-  • $S_{\text{uncached}}$: average number of tokens in prefill phase not hit by prefill cache (requires recomputing $kv_{\text{cache}}$)
-  • attn_heads: number of attention heads
-  • kv_heads: number of separate key/value heads in GQA
-  • head_size: dimension of each head
-  • inter_size: intermediate dimension of FFN layer
-  • vocab_size: vocabulary size
-  • $L$: number of model layers
-Following [122], we consider only the floating-point operations (FLOPs) from matrix multiplications (GEMMs). For the Qwen model:
+  • $b$: the batch size
+  • $h$: the hidden size
+  • $s_{\text{t}}$: the total average number of prompt tokens
+  • $s_{\text{u}}$: the average number of uncached tokens in prefill phase
+  • $n_{\text{a}}$: the number of attention heads
+  • $n_{\text{kv}}$: the number of separate key/value heads in GQA
+  • $d_{\text{h}}$: the dimension of each head
+  • $d_{\text{i}}$: the intermediate dimension of FFN layer
+  • $v$: the vocabulary size
+  • $l$: the number of model layers
+According to [84], we calculate the floating-point operations (FLOPs) of the Qwen2.5 models, only considering matrix multiplications (GEMMs) without norms and activations:
   • Attention block operations:
-    • Query/Key/Value transformations (GQA):
-[…]
-    • Attention matrix computation and attention over values:
-[…]
-    • Post-attention linear projection:
-[…]
+    • Query/Key/Value transformations (GQA): $2bs_{\text{u}}hd_{\text{h}}(n_{\text{a}}+2n_{\text{kv}})$
+    • Attention matrix computation and attention over values: $2bs_{\text{t}}s_{\text{u}}n_{\text{a}}d_{\text{h}}$
+    • Post-attention linear projection: $2bs_{\text{u}}hn_{\text{a}}d_{\text{h}}$
   • Feed-forward network block operations:
-    • Up/Gate/Down projections (each):
-[…]
-[…]
+    • Up/Gate/Down projections (each): $2bs_{\text{u}}hd_{\text{i}}$
+  • LM head operations: $2bhv$
 Thus, the total FLOPs for one inference with prefill cache is:
-[…]
-[…]
-[…]
-[…]
+$\text{FLOPs}_{\ \text{total}}$ $=l\Bigg{[}4bs_{\text{u}}hd_{\text{h}}\left(n_{\text{a}}+n_{\text{kv}}\right)+2bs_{\text{u}}s_{\text{t}}d_{\text{h}}n_{\text{a}}+6bs_{\text{u}}hd_{\text{i}}\Bigg{]}+2bhv$
 Workload and Computational Analysis
-Our experiments workload consists of prompts with an average length of 1792 tokens.
-By leveraging prefix caching to exploit shared prompt prefixes during inference, we reduce the average number of uncached tokens in the prefill phase to 128, while generating an average of 9 tokens per request.
-For FLOPs calculations, we account for $S_{\text{uncached}}=128$ tokens in the prefill phase and $S_{\text{uncached}}=1$ token per decoding step, with detailed results presented in Table 2.
+We conduct experiments on our datasets, in which prompts have an average length of 1792 tokens.
+By leveraging prefix caching to exploit shared prompt prefixes during inference, we reduce the average number of uncached tokens in the prefill phase to 128 and generate an average of 9 tokens per request.
+We calculate FLOPs under $s_{\text{u}}=128$ tokens in the prefill phase and under $s_{\text{u}}=1$ token for each step in the decoding phase, with detailed results presented in Table 2.
+Appendix K Extensive Attempts on Generalization Enhancement via RL
+Table 11: Full results during the RL exploration
+Model AR
+AR
+(Out-of-Domain)
+Crowdsourced Labelers 77.85% -
+BERT-based System 81.43% -
+Qwen2.5-72B (ICL) 94.74% 94.63%
+Qwen2.5-7B (RFT) 95.29% 90.86%
+Qwen2.5-3B (RFT) 95.30% 91.02%
+Qwen2.5-0.5B (RFT) 94.66% 90.62%
+Qwen2.5-0.5B (KD) 94.66% 90.62%
+Qwen2.5-7B (RL) 95.36% 94.80%
+Qwen2.5-3B (RL) 94.68% 92.94%
+Qwen2.5-0.5B (RL) 90.76% 85.21%
+Qwen2.5-0.5B (KD*) 88.49% 84.14%
+Qwen2.5-0.5B (KD*+RL) 95.55% 91.68%
+While attempting to apply reinforcement learning (RL), we conduct extensive experiments, with results shown in Appendix K.
+Initially, we find that directly training a 7B model with RL yields excellent results, surpassing the LLM Prototype and RFT-trained models.
+Subsequently, we attempt to train smaller models (3B and 0.5B) and discover that RL is ineffective on the 0.5B model, because the model tends to generate shorter or even meaningless Chain-of-Thought (CoT), leading to degraded output diversity as well as that without CoT.
+To address the issue above, we explore another approach, transferring out-of-domain capabilities into the 0.5B model via Knowledge Distillation (KD).
+However, we observe that the model struggles with format errors due to underfitting.
+Ultimately, after applying RL training on this model, we notice an impressive improvement on model performance, with both an absolute 1% AR increase on in-domain/out-of-domain datasets compared to the 0.5B RFT-trained model.

### v3

@@
 Abstract
-In recent years, Large Language Models (LLMs) have significantly advanced artificial intelligence by optimizing traditional Natural Language Processing (NLP) pipelines, improving performance and generalization.
-This has spurred their integration into various systems.
-Many NLP systems, including ours, employ a "one-stage" pipeline directly incorporating LLMs.
-While effective, this approach incurs substantial costs and latency due to the need for large model parameters to achieve satisfactory outcomes.
-This paper introduces a three-stage cost-efficient end-to-end LLM deployment pipeline—including prototyping, knowledge transfer, and model compression-to tackle the cost-performance dilemma in LLM-based frameworks.
-Our approach yields a super tiny model optimized for cost and performance in online systems, simplifying the system architecture.
-Initially, by transforming complex tasks into a function call-based LLM-driven pipeline, an optimal performance prototype system is constructed to produce high-quality data as a teacher model.
-The second stage combines techniques like rejection fine-tuning, reinforcement learning, and knowledge distillation to transfer knowledge to a smaller 0.5B student model, delivering effective performance at minimal cost.
-The final stage applies quantization and pruning to extremely compress models to 0.4B, achieving ultra-low latency and cost.
-The framework’s modular design and cross-domain capabilities suggest potential applicability in other NLP areas.
-* Equal Contribution Corresponding Author
+Large Language Models (LLMs) have significantly advanced artificial intelligence by optimizing traditional Natural Language Processing (NLP) workflows, facilitating their integration into various systems.
+Many such NLP systems, including ours, directly incorporate LLMs.
+However, this approach either results in expensive costs or yields suboptimal performance after fine-tuning.
+In this paper, we introduce a three-stage cost-efficient end-to-end LLM deployment pipeline, comprising prototyping, knowledge transfer, and model compression, to effectively tackle the cost-performance dilemma in LLM-based frameworks.
+Its high cost-efficiency is manifested not only in simplifying system complexity and producing super-tiny online models with enhanced performance and reduced costs in the results, but also in addressing development cycle constraints, the lack of extensive high-quality data, and limited computational resources during the project development process.
+In the first stage, we construct an optimal performance prototype system by transforming complex tasks into a function call-based LLM-driven pipeline, which serves as a teacher model to generate high-quality data.
+In the second stage, we combine techniques like rejection sampling fine-tuning, reinforcement learning, and knowledge distillation to transfer knowledge to 0.5B student models, delivering effective performance at minimal cost.
+In the final stage, we further compress models to 0.4B via quantization and pruning, achieving ultra-low latency and cost.
+Extensive experimental results and the framework’s modular design suggest cross-domain capabilities and potential applicability in other NLP areas.
+[…]
+[…]
 1 Introduction
-Figure 1: From the traditional BERT-based system to the LLM-based system.
-In recent years, Large Language Models (LLMs) have made significant strides [1; 2; 3] in enhancing artificial intelligence by addressing shortcomings of traditional pipelines — including manual annotation, suboptimal performance, and limited generalization and adaptability.
+In recent years, Large Language Models (LLMs) have made significant strides [1, 2, 3] in enhancing artificial intelligence by addressing shortcomings of traditional pipelines — including manual annotation, suboptimal performance, and limited generalization and adaptability.
 Consequently, LLM-based pipelines have stimulated substantial demand for integration into various Natural Language Processing (NLP) systems.
-We find that numerous NLP systems, including our own, have directly adopted LLM into the pipeline, often referred to as the "one stage" pipeline.
-Despite the breakthroughs facilitated by this pipeline, it also incurs substantial cost increases [4] and high inference latency [5].
+We find that numerous such NLP systems, including our own, have directly utilized APIs of closed-source LLMs or fine-tuned open-source LLMs, which can be referred to as the "one-stage" pipeline.
+Although this pipeline offers implementation simplicity, it suffers from substantial costs [4] and high inference latency [5] or suboptimal performance, which limit the broader applicability of LLMs and impact their commercial viability.
 Within our NLP system, cost-performance is an especially critical metric, a sentiment likely shared by others in the field [6].
-These significant costs limit the broader applicability of LLMs and impact their commercial viability.
-Our analysis points out that it is precisely the simplistic training method of the one-stage pipeline, which necessitates larger model parameters to achieve satisfactory results, that leads to significant costs.
-In this paper, we introduce a three-stage performant and cost-efficient end-to-end LLM deployment pipeline, consisting of prototyping, knowledge transfer, and model compression, to address the pervasive "cost-performance" dilemma in the practical deployment of LLM-based frameworks.
-This approach ultimately yields a super tiny model, which is deployed in our online system, achieving optimal cost-performance metrics while substantially simplifying system architecture.
-The first stage involves constructing the prototype system, where traditional complex tasks are converted into function call-based LLM-driven workflows [7] through prompt engineering, resulting in an optimal performance prototype compared to traditional pipeline.
+Moreover, our analysis indicates that this pipeline struggles to balance the cost and the performance due to insufficient high-quality data and monotonous optimization techniques.
+In this paper, we introduce a three-stage cost-efficient end-to-end LLM deployment pipeline, as shown in Figure 1, consisting of prototyping, knowledge transfer, and model compression, to address the pervasive "cost-performance" dilemma in the practical deployment of LLM-based frameworks.
+The high cost-efficiency of this approach is demonstrated by both the outcomes and the process.
+In terms of outcomes, the approach simplifies system complexity and produces super-tiny models with optimal performance and reduced inference costs and latency.
+Regarding the process, the approach addresses challenges such as constrained development cycles, the scarcity of substantial high-quality data, and limited computational resources.
+In our proposed pipeline, the first stage involves constructing the prototype system, where traditional complex tasks are converted into function call-based LLM-driven workflows [7] through prompt engineering, resulting in an optimal performance prototype compared to traditional pipeline.
 This prototype serves as a teacher model in the second stage, efficiently generating vast quantities of high-quality data.
-During the knowledge transfer phase, leveraging the optimal prototype from the first stage, we employ techniques including rejection fine-tuning (RFT), reinforcement learning (RL) [8; 9; 10], and multi-strategy knowledge distillation (KD) targeting the Qwen 2.5 series model [11], effectively transferring knowledge to a 0.5B student model.
+In the second stage, leveraging the optimal prototype from the first stage, we employ techniques including Rejection sampling Fine-Tuning (RFT), Reinforcement Learning (RL) [8, 9, 10], and multi-strategy knowledge distillation (KD) on the Qwen 2.5 series model [11].
+Our hybrid strategy effectively transfers generalization capabilities and the abilities of generating high-quality CoT to a 0.5B student model.
 The derived student model is considerably smaller than the prototype model while achieving nearly comparable performance at minimal cost.
-In the third stage, we utilize quantization and pruning techniques to further compress the student model developed in the second phase, resulting in a remarkably compact model of 0.4 billion parameters (as shown in Figure 1).
-Ultimately, this super tiny model is deployed in our online system.
+In the third stage, we utilize quantization and pruning techniques to further compress the student model developed in the second phase, resulting in a remarkably compact 0.4B model.
+Ultimately, this super-tiny model is deployed in our online system.
 This model achieves hundreds-fold compression and extremely low latency and cost in exchange for a trivial and acceptable loss in performance.
 The modular design and superior cross-domain performance of the framework suggest its potential applicability in other NLP domains.
-The major contributions of this paper are as follows:
-  • The first validated three-stage cost-efficient end-to-end LLM pipeline with simplifying original system complexity achieves SOTA results and resolves the "cost-performance" dilemma, realizing a 180-fold compression with nearly consistent AR compared to 72B LLMs and up to an absolute 14% AR increase compared to traditional BERT-based pipelines.
-  • We propose the first hybrid strategy that significantly enhances the generalization of super tiny models on domain-specific tasks through combining RL and KD.
-This approach effectively transfers the capabilities of LLMs to smaller models.
+Our major contributions are as follows:
+  • The first validated three-stage cost-efficient end-to-end LLM pipeline, simplifying system complexity, achieves SOTA results and resolves the "cost-performance" dilemma.
+It realizes 180 times compression with nearly consistent performance compared to original LLMs and up to an absolute 14% performance increase over traditional BERT-based systems.
+  • We propose the first hybrid knowledge transfer method by combining RL and KD.
+It effectively transfers the capabilities of LLMs to super-tiny models and achieves performance comparable and even superior to certain larger models on domain-specific tasks.
   • The proposed framework offers advantages such as low training, inference, and deployment costs, minimal project development expenses with fast cycles, and a strong potential for cross-domain transferability.
 It provides inspiration for developing standardized cross-domain solutions in both academia and industry, facilitating the widespread LLMs adoption.
+Figure 1: Workflows of prototyping, knowledge transfer and compression.
 2 Related Works
@@
 Query-to-API reformulation via rejection sampling can also optimize API retrieval, reducing retrieval latency while improving accuracy [15].
-Nowadays, LLMs can be used to evaluate model-generated content [16; 17].
+Nowadays, LLMs can be used to evaluate model-generated content [16, 17].
 CriticGPT [18] indicates that LLMs can build a cost-effective and high-precision judge, surpassing humans in detecting subtle errors.
@@
 Knowledge distillation (KD) [19] transfers knowledge from a teacher model to a smaller student model by simulating probability distribution.
-Common approaches include Logits-based[20; 21], Feature-based [22] and sequence-level (Seq-KD) [23] distillation.
-Initially, Logits-based KD uses the forward KL divergence (FKL) [24; 25] as the loss function.
-Some KD methods prefer Reverse KL divergence (RKL) [26; 27; 28; 21].
+Common approaches include Logits-based[20, 21], Feature-based [22] and sequence-level (Seq-KD) [23] distillation.
+Initially, Logits-based KD uses the forward KL divergence (FKL) [24, 25] as the loss function.
+Some KD methods prefer Reverse KL divergence (RKL) [26, 27, 28, 21].
 Recently, Adaptive Kullback-Leiber (AKL) divergence [29] achieves a balance between mimicking distribution and avoiding overfitting.
 LLM Pruning & Quantization
-Efficient LLM deployment primarily relies on pruning[30; 31] and quantization[32; 33; 34].
-Structured pruning[35] employs layer removal[36; 37] or dimensionality reduction[31; 37; 38] using gradient-based[31; 38] or activation-aware[37] importance criteria, implemented via learnable masks[38; 39] or pruning-retraining pipelines[36; 31; 37].
-Weight-only LLM quantization methods[32; 33; 40] improve memory efficiency while joint weight-activation schemes[34; 41] and their combination [42] accelerate computation.
+Efficient LLM deployment primarily relies on pruning[30, 31] and quantization[32, 33, 34].
+Structured pruning[35] employs layer removal[36, 37] or dimensionality reduction[31, 37, 38] using gradient-based[31, 38] or activation-aware[37] importance criteria, implemented via learnable masks[38, 39] or pruning-retraining pipelines[36, 31, 37].
+Weight-only LLM quantization methods[32, 33, 40] improve memory efficiency while joint weight-activation schemes[34, 41] and their combination [42] accelerate computation.
 Reinforcement Learning
-Recent works [8; 9; 10] focus on leveraging Reinforcement Learning (RL) to enhance model reasoning capabilities.
-Notable reasoning models such as DeepSeek-R1 [43], QWQ-32B [44], and OpenAI o1 [45] have emerged, employing RL algorithms like Proximal Policy Optimization (PPO) [46], Group Relative Policy Optimization (GRPO) [47], and others [48; 49; 50].
-For general-purpose LLMs, preference optimization algorithms [51; 52; 53; 54] are widely adopted.
-Additionally, self-improving pipelines [55; 56; 57] have been proposed to iteratively refine model performance.
+Recent works [8, 9, 10] focus on leveraging Reinforcement Learning (RL) to enhance model reasoning capabilities.
+Notable reasoning models such as DeepSeek-R1 [43], QWQ-32B [44], and OpenAI o1 [45] have emerged, employing RL algorithms like Proximal Policy Optimization (PPO) [46], Group Relative Policy Optimization (GRPO) [47], and others [48, 49, 50].
+For general-purpose LLMs, preference optimization algorithms [51, 52, 53, 54] are widely adopted.
+Additionally, self-improving pipelines [55, 56, 57] have been proposed to iteratively refine model performance.
 We primarily utilize RL to enhance the model’s generalization.
@@
 3 Methodology
-Figure 2: Overview of Prototype, Production and Application.
 Recent advancements in LLMs have facilitated significant progress in NLP tasks such as classification and summarization.
-We harness these capabilities to reconstruct the traditional dialogue system through a cost-efficient end-to-end LLM pipeline, which includes prototyping, knowledge transfer, and model compression, as illustrated in 2(a) and 2(b).
+We harness these capabilities to reconstruct the traditional dialogue system through a cost-efficient end-to-end LLM pipeline, which includes prototyping, knowledge transfer, and model compression, as illustrated in 1(a) and 1(b).
 Initially, we develop the prototype system by leveraging the exceptional performance of contemporary LLMs.
@@
 Finally, we achieve additional reduction in model size through further model compression techniques.
-Ultimately, our objective is to deliver intelligent user experiences at minimal costs, as demonstrated in 2(c).
+Ultimately, our objective is to deliver intelligent user experiences at minimal costs, as demonstrated in 1(c).
 3.1 Prototyping
-Table 1: Complexity statistics of Our dataset vs Open Source datasets. More details in Appendix A.
-Dataset Test Set Size Intents Slots TTR Entropy
-TinyAgent[14] 1023 17 48 0.0087 7.8220
-Snips[58] 893 7 72 0.0207 7.6997
-ATIS[59] 700 21 120 0.0111 7.4148
-Our dataset 11922 74 225 0.1410 9.4250
 LLM Prototype
-Traditional dialogue systems usually employ the Intent Classification and Slot Filling (ICSF) framework, often utilizing BERT[60; 61].
-However, a single model often struggles to handle complex real-world issues, as shown in the Table 1.
-Consequently, a Domain Classifier (DC) is frequently used to first categorize queries into domains, enabling domain-specific ICSF parsing.
+Traditional dialogue systems usually employ the Intent Classification and Slot Filling (ICSF) framework, often utilizing BERT[58, 59].
+Given the limitations of a single model in handling complex real-world issues, a Domain Classifier (DC) is often utilized for initial query categorization to facilitate domain-specific ICSF parsing.
 Although the hierarchical DC-ICSF architecture improves accuracy, the system complexity increases with feature updates.
@@
 Following this approach, we design a prototype system based on Qwen2.5-72B-Instruct[11].
-As illustrated in 2(a), the prototype system accepts user queries and contextual information (including chat history and temporal status information) as input, searches relevant APIs from the API pool, then outputs function names and parameters to call.
+As illustrated in 1(a), the prototype system accepts user queries and contextual information (including chat history and temporal status information) as input, searches relevant APIs from the API pool, then outputs function names and parameters to call.
 The traditional system comprises one DC module and multiple ICSF models (based on BERT or regular expressions), each requiring separate maintenance and frequent DC updates.
-In contrast, the LLM-based system, equipped with function call capability, processes all tasks through a unified model, as illustrated in Figure 1, thereby reducing system complexity and enhancing operational efficiency.
+In contrast, the LLM-based system, equipped with function call capability, processes all tasks through a unified model, thereby reducing system complexity and enhancing operational efficiency.
 To enhance temporal awareness and mitigate hallucinations, we enrich the input with contextual knowledge via RAG [60], retrieving explanations and extensions of entity-related terms from a curated knowledge base.
@@
 Then, the LLM selects optimal functions/parameters from candidate APIs.
-If no suitable function exists, the system returns "No appropriate function"; in the case of ambiguous queries, "Request unclear" are generated.
-For backward compatibility, we design an Adapter to map functions/parameters to legacy domain-intent-slot structures.
+For backward compatibility, we map functions/parameters to legacy domain-intent-slot structures.
 Judge
-The traditional evaluation scheme for ICSF relies on the test set that requires full alignment of Intent and Slots.
-However, in real-world scenarios, there exist multiple solutions for function and slot selection.
-For example: (1) "Slightly increase the volume" could be resolved using either volume setting or volume increasing APIs; (2) "11:00 pm" and "23:00" semantically represent the same time.
-Additionally, due to crowdsourced annotators’ limited understanding of specific business logic, the manual labeling accuracy remains suboptimal.
-As shown in 2(c), the crowdsourced labeler’s annotation accuracy has become the system bottleneck.
+Traditional ICSF evaluation relies on test sets demanding strict intent-slot alignment, contrasting with real-world variability where utterances permit multiple valid interpretations.
+For example, "Slightly increase the volume" allows alternative API calls, and "11:00 pm"/"23:00" are semantically equivalent.
+This inherent ambiguity and complex business logic significantly impede consistent human annotation for accurate ground truth.
 To address issues above, we introduce LLM-as-a-Judge to evaluate prototype accuracy.
@@
 Based on these findings, we conclude that LLM-as-a-Judge is the optimal evaluation approach for this task.
-As demonstrated in Figure 4, we reconstruct the dialogue pipeline from BERT-based DC-ICSF architecture to LLM-based Function Call architecture.
+We reconstruct the dialogue pipeline from BERT-based DC-ICSF architecture to LLM-based Function Call architecture.
 Evaluation through LLM-as-a-Judge demonstrates that the new architecture improves accuracy from 81.43% to 94.74% compared to the original one.
@@
 For a dataset $D$ containing prompt component $x$ and response component $y$, we define $D(x,y)=\{x_{i},y_{i}\}_{i=1}^{N}$, where $N$ denotes the dataset size.
-To minimize the Cross Entropy (CE) loss, we first utilize the SFT optimization, and then employ further optimization Rejection sampling Fine-Tuning (RFT).
-We use LLM judge introduced in Section 3.1 to evaluate training data generated by the LLM prototype, filter out low-quality data, and generate the final training set.
+SFT employs Cross Entropy (CE) loss [63] as the optimization objective.
+Building on this, we apply Rejection sampling Fine-Tuning (RFT), utilizing an LLM judge, as introduced in Section 3.1, to assess training data generated by the LLM prototype, filter out low-quality data, and construct the final training dataset.
 The actual training data is formulated as:
@@
 […]
-Reinforcement Learning
-Since OpenAI introduced Reinforcement Learning from Human Feedback (RLHF) [65], it has become an essential component for aligning LLMs, as evidenced in major technical reports [66; 67; 68; 11; 69].
-DeepSeek-R1 [43] proposes a cost-effective RL training method from scratch which combined rule-based ORM and GRPO [47].
-Subsequent works like Logic-RL [8; 9; 10] applies this efficient approach across various tasks, enabling the model to automatically generate Chain of Thought (CoT) during training, thereby enhancing performance [70].
-Based on our observation that model generalization degrades after RFT, we attempt to employ RL to improve model performance on out-of-domain datasets.
-Following Logic-RL [9], we refine the reward design for our specific task (see Appendix B) and conduct extensive attempts to integrate knowledge distillation (see Appendix K).
-Distillation
-Using the Function Call-related data from large models and the RFT methods, we transfer capabilities from large to small models through training.
-To accelerate model training and inference, we implement prompt compression during training by removing unnecessary content from input-output pairs, including format specifications in inputs and Chain-of-Thought (CoT) rationales in outputs.
-Please refer to Appendix C for details.
-Knowledge distillation [19] is an effective method for transferring knowledge from large to small models, requiring fewer data and exhibiting lower overfitting risks compared to SFT.
-We adopt the logits-based Adaptive Kullback-Leibler (AKL) distillation [29], with the optimization objective:
+Knowledge Distillation
+Knowledge Distillation (KD) [19] offers an alternative to RFT for transferring capabilities from large to small models, typically requiring fewer data samples and exhibiting lower overfitting risks.
+We adopt the logits-based Adaptive Kullback-Leibler (AKL) distillation [29]:
 $L_{AKL}=\mathbb{E}_{(x,y)\sim D}[AKL(\pi_{\phi}(y\lvert x),\pi_{\theta}(y\lvert x))]$ (3)
 $AKL(p,q)=\alpha_{head}KL(p,q)+(1-\alpha_{head})KL(q,p)$ (4)
-$KL(p,q)=\sum_{k}{p_{i}}\log\frac{p_{i}}{q_{i}}$ (5)
+$KL(p,q)=\sum_{t}{p_{t}}\log\frac{p_{t}}{q_{t}}$ (5)
 where $\pi_{\theta}$ denotes the trainable student model, $\pi_{\phi}$ represents the teacher model, and $\alpha_{head}$ is the head ratio in AKL.
@@
 We make optimizations on original methods: (1) Teacher Calibration: Following [64], we calibrate the teacher model with task-specific data to achieve lower student model loss; (2) Logits Pre-storage: For reducing training time, we pre-store teacher logits to enable multi-student distillation from a single cached logits set; (3) Top-k Pruning: We trim low-probability logits through top-k pruning, mitigating storage pressure and accelerating training speed.
+Hybrid Knowledge Transfer
+Figure 2: The workflow of the hybrid knowledge transfer strategy.
+Following the standard knowledge transfer strategy, we employ either RFT or KD combined with prompt compression, which removes unnecessary format specifications and Chain-of-Thought (CoT) rationales (see Appendix B), to create a fast but domain-specific model.
+However, our experiments reveal limitations of this conventional approach.
+Utilizing RFT or KD alone can improve inference efficiency but fails to enhance out-of-domain generalization.
+According to Logic-RL [9] with task-specific reward adjustment (see Appendix A), we employ Reinforcement Learning (RL) to improve generalization.
+Additional experiments indicate that employing RL alone generally leads to varying effectiveness across different model sizes: RL training yields exceptional results on small models (7B), surpassing both the LLM prototype and RFT-trained models.
+On the other hand, when applied to super-tiny models (0.5B), RL becomes ineffective as these models tend to generate shorter or meaningless CoT responses, leading to degraded output diversity.
+These findings motivate us to propose a hybrid knowledge transfer strategy, as illustrated in Figure 2.
+This hybrid strategy first employs RL to train models that are as small as possible while ensuring robust generalization and the ability to generate high-quality CoT, exemplified by the 7B RL-trained model in our experiments.
+Subsequently, KD is used to transfer these generalization capabilities and CoT generation abilities to super-tiny models.
+With CoT generation abilities, super-tiny models can be trained through a second round of RL to further enhance generalization.
+Experimental results demonstrate that super-tiny models exhibit impressive improvement in both in-domain and out-of-domain performance after the second RL training.
+Detailed performance comparison of different transfer strategies shows in Appendix J.
 3.3 Model Compression
@@
 For depth pruning, we compute layer importance scores and remove the $N$ least important layers based on the target pruning ratio.
-For Width Pruning, we conduct a model architecture search using PPL-based criteria and prune channels according to group-wise importance rankings, with full implementation details in Section I.1; (2) Recovery training: The first step is General Instruction Fine-tuning via Knowledge Distillation.
+For width pruning, we conduct a model architecture search using PPL-based criteria and prune channels according to group-wise importance rankings, with full implementation details in Appendix H; (2) Recovery training: The first step is General Instruction Fine-Tuning via Knowledge Distillation.
 Adopting Minitron’s approach, we leverage logits from a larger general instruction teacher model for recovery training.
@@
 In our experiments, we utilize a large-scale dataset built from years of team experience and practice accumulated in the field.
-We reconstruct the intent classification and semantic slot framework within the real-world system into an API document specification, creating a function pool comprising 74 functional modules and 225 parameters.
+We reconstruct the intent classification and semantic slot framework within the real-world system into an API document specification, creating a function pool comprising 74 functional modules with 225 arguments.
 We sample 200k data points from a single business scenario as the training set and synthesize labels based on the LLM prototype outlined in Section 3.1.
@@
 Additionally, we sample 7,867 data points using the same methodology from another business scenario, without some callable functions in the training set, as the out-of-domain test set.
-Compared to open-source datasets, our self-constructed dataset is more complex and closely aligned with real-world scenarios (see Table 1).
 Metrics
@@
 We also apply KD from 7B RL-trained models to 0.5B models, followed by another RL phase (denoted as KD*+RL).
-Finally, we use model quantization via W8A16/W4A16/W8A8 strategies on the KD-trained 0.5B model, detailed in Appendix E.
+Finally, we use model quantization via W8A16/W4A16/W8A8 strategies on the KD-trained 0.5B model, detailed in Appendix D.
 The training process is cost-efficient, utilizing only 8 NVIDIA H20 GPUs.
-Computing resources are listed in Appendix F.
+Computing resources are listed in Appendix E.
 Evaluation settings
-We do evaluation by comparing our models with crowdsourced labelers and the BERT-based system consisting of two models based on BERT-large (330M)[75].
+We do evaluation by comparing our models with the BERT-based system consisting of two models based on BERT-large (0.33B)[68].
 To evaluate the effectiveness of our method, we invite expert annotators alongside the LLM-as-a-Judge-based metric AR (see Equation 6) for performance assessment.
-We also implement a retrieval module to select top 10 functions from the function pool, aiming to accelerate both training and inference phases, as detailed in Appendix D.
-In real-world scenarios, we deploy models on NVIDIA L20 GPU and perform effectiveness evaluation via TensorRT-LLM[74], with specific configuration details provided in Appendix H.
+We also implement a retrieval module to select top 10 functions from the function pool, aiming to accelerate both training and inference phases, as detailed in Appendix C.
+To further verify the effectiveness of our proposed approach, we train models on the xlam-function-calling-60k dataset generated by APIGen[69] and conduct validation using the Berkeley Function-Calling Leaderboard (BFCL)[70].
+Details can be found in Appendix K.
+In real-world scenarios, we deploy models on NVIDIA L20 GPU and perform effectiveness evaluation via TensorRT-LLM[67], with specific configuration details provided in Appendix G.
 During the comparison between the trained models and the BERT-based system, we adopt queries per second (QPS) as the primary metric for system throughput, response time (RT) for end-to-end latency, and concurrent clients (CC) to quantify high-concurrency performance.
 4.2 Experimental Results
-Table 2: Main results of the models trained on Qwen2.5(See Appendix J for FLOPs analysis).
-Model / System AR GFLOPs
-Crowdsourced Labelers 77.85% -
+Table 1: Main results of the models trained on Qwen2.5 (See Appendix I for FLOPs analysis).
+Model / System AR FLOPs (G)
 BERT-based system 81.43% 38.65
@@
 Qwen2.5-0.5B-W8A8 (KD) 94.83% -
-Qwen2.5-0.4B-Width (KD) 94.17% 33.10
-Qwen2.5-0.4B-Depth (KD) 94.20% 40.83
+Qwen2.5-0.4B-Width (KD) 94.17% 40.83
+Qwen2.5-0.4B-Depth (KD) 94.20% 33.10
 Figure 3: Cost and performance results. FLOPs reduced by 10x with nearly lossless AR.
-Table 3: RFT vs RL on in-domain data and out-domain data.
+Table 2: In-domain and out-of-domain performances of different knowledge transfer strategies.
 Model AR
@@
 Qwen2.5-7B (RL) 95.36% 94.80%
-Qwen2.5-0.5B (KD*+RL) 95.55% 91.68%
-Figure 4: Comparison over different models by professional annotators
+Qwen2.5-0.5B (KD*+RL) 95.99% 92.34%
+Figure 4: Comparison of annotation results between the LLM prototype and the BERT-based system determined by professional annotators.
 LLM Prototype Preferred by Both Human and LLM-as-a-Judge
-As illustrated in Figure 4, annotations provided by professional annotators reveal that the LLM prototype (72B ICL) achieves a win+tie rate of 95.59% over the BERT-based system and 92.98% over crowdsourced labelers.
-This demonstrates the superior performance of the LLM prototype.
-Through the evaluation via LLM-as-a-Judge, we find that the LLM Prototype surpasses the BERT-based System by 13.31% and crowdsourced annotations by 16.89%, as shown in Table 2.
+As illustrated in Figure 4, annotations provided by professional annotators reveal that the LLM prototype (72B ICL) achieves a superior performance over the BERT-based system, with a win+tie rate of 95.59%.
+Through the evaluation via LLM-as-a-Judge, we find that the LLM Prototype surpasses the BERT-based system by 13.31%, as shown in Table 1.
 This conclusion is nearly consistent with that presented by professional annotators, demonstrating the effectiveness of the LLM judge.
@@
 Conversely, 7B RL-trained models achieve superior performance compared to other models on both in-domain and out-of-domain datasets, indicating better generalization of RL training methodology.
-Notably, even with models as small as 0.5B, employing a hybrid training methodology with KD+RL can yield superior in-domain results and an improvement of 1% in out-of-domain datasets compared to the RFT method.
-For more detailed exploration of the RL results, please refer to Appendix K.
-Tiny LLMs as a Cost-Efficient Upgrade to BERT-based System
-As shown in Table 2 and Table 4, quantization minimally impacts accuracy, with less than a 0.1% AR loss, while significantly enhancing inference performance, achieving up to 1.29x higher QPS and 0.76x lower RT in the optimal scenario.
+Notably, even with models as small as 0.5B, employing a hybrid training methodology with KD+RL can yield superior in-domain results and an improvement of 1.7% in out-of-domain datasets compared to the RFT method.
+For more detailed exploration of the RL results, please refer to Appendix J.
+Tiny LLMs as a Cost-Efficient Upgrade of BERT-based System
+As shown in Table 1, Table 3 and Figure 5, quantization minimally impacts accuracy, with less than a 0.1% AR loss, while significantly enhancing inference performance, achieving up to 1.29x higher QPS and 0.76x lower RT in the optimal scenario.
 Pruning incurs a slightly higher 0.2% accuracy loss but consistently enhances performance by 20%.
@@
 Although the cost is roughly twice that of the BERT-based system, the AR improves by up to absolute 14%.
-Table 4: Inference Performance under different model compression techniques.
+Table 3: Comparison of QPS metrics using different quantization and pruning strategies.
 Model QPS
@@
 Qwen2.5-7B 4.43 14.77 17.71 19.04 19.79 20.06
-Figure 5: Normalized RT ratio based on FP16 models
+Figure 5: Comparison of RT metrics using different quantization and pruning strategies.
 5 Analysis
@@
 Depth pruning achieves 20% inference acceleration with lossless AR at same compression ratios (Table 1), stemming from the nature of depth pruning, removing entire transformer layers and their associated attention computations.
-Although width pruning exhibits slightly better effects in general instruction tasks after training (Table 9), depth pruning presents optimal efficiency facing long prompts, with width pruning focusing on matrix optimizations (Appendix J).
+Although width pruning exhibits slightly better effects in general instruction tasks after training (Table 8), depth pruning presents optimal efficiency facing long prompts, with width pruning focusing on matrix optimizations (Appendix I).
 Regarding quantization, the W4A16 strategy achieves impressive performance in single-thread scenarios, while FP8 outperforms with lossless accuracy under high concurrency conditions as the superior paradigm for throughput deployments.
 6 Conclusion
-Based on our comprehensive study, several key insights emerge regarding the development and deployment of effective deep learning models in practical scenarios.
-Tiny Models as a Practical Low-Cost Alternative
-Our experiments indicate that end-to-end optimized tiny models can achieve performance levels comparable to their larger counterparts, offering resource efficiency without compromising accuracy.
+Based on our comprehensive study, several key insights emerge regarding the development and deployment of effective LLMs in practical scenarios.
+A Scalable Practical Alternative for Cost-Efficient LLMs
+Our proposed pipeline presents a systematic framework designed to achieve high performance while maintaining efficient costs in both development and deployment.
+Our end-to-end optimized tiny models also outperform their larger counterparts, offering resource efficiency without compromising accuracy.
 This provides an alternative pathway for applying LLMs in practical tasks under low-cost scenarios.
+We believe this approach can be effectively adapted to other domains, thereby potentially advancing research and facilitating the widespread deployment of LLMs.
+LLM Capability Preservation via a Hybrid Strategy
+Our hybrid strategy demonstrates strong effectiveness and practicality in preserving LLMs’ capabilities during knowledge transfer.
+It enhances generalization and high-quality CoT generation abilities via RL, safely retains these capabilities in smaller models via KD, and further improves the in-domain and out-of-domain performance of super-tiny models via another RL.
+This methodology judiciously and scientifically avoids irreversible loss of capabilities that might occur when directly deriving super-tiny models via RL.
+It progressively retains LLMs’ capabilities, transferring them step by step to super-tiny models.
 Balanced Focus Beyond AGI
@@
 It advocates for a balanced research agenda that equally emphasizes foundational advancements and domain-specific applications.
-Scalable Cost-Efficient Three-Stage Pipeline
-The proposed pipeline-comprising prototyping, knowledge transfer, and model compression-presents a systematic framework designed to achieve high performance while maintaining low costs in both training and inference.
-We believe that this approach can be effectively adapted to other domains, thereby potentially advancing research in those areas and facilitating the widespread deployment of LLMs.
 7 Limitation & Future Works
-Our pipeline has presented optimal cost-performance results, and our super tiny models have exhibited remarkable capabilities, featuring low latency, cost efficiency, cloud independence, and robust privacy.
+Our pipeline has presented optimal cost-performance results, and our super-tiny models have exhibited remarkable capabilities, featuring low latency, cost efficiency, cloud independence, and robust privacy.
 Further compression may offer limited marginal benefits.
-Thus, we are more focused on enhancing such models’ capabilities and expand their broader applications such as mobile devices:
-Hybrid Methodology for LLMs Capabilities Transfer
-Based on our existing experiments, the mixer of reinforcement learning and knowledge distillation has effectively enhanced the generalization capabilities of super tiny models.
-This implies a potentially powerful hybrid methodology that allows super tiny models to nearly fully retain the capabilities of LLMs in domain-specific tasks.
-In future research, we will build upon our previous work to conduct deeper studies and theoretical analyses of this hybrid approach in transferring LLMs capabilities to such super tiny models.
-On-device Deployment Optimization
-Although we achieve 8-bit compression and online deployment, real-world edge devices face heterogeneous hardware constraints such as memory fragmentation and NPU compatibility.
-In the future, for on-device deployment of super tiny models, we will develop hardware-aware compression algorithms and adaptive runtime kernels to optimize latency-energy tradeoffs across diverse IoT ecosystems, and explore more resource-efficient deployment techniques for edge devices, such as sparse computation, edge-centric quantization, and federated learning.
-Appendix A Dataset Analysis
-Type-Token Rate (TTR), as a fundamental measure of lexical diversity, reflects the local variability of vocabulary, but it is sensitive to text length. The calculation formula is:
-$TTR=\frac{N_{types}}{N_{tokens}}$ (7)
-where $N_{tokens}$ denotes the total number of 2-gram patterns in the text, and $N_{types}$ represents the number of distinct 2-gram patterns.
-Entropy, as an information-theoretic metric, quantifies the uncertainty within binary grammatical sequences based on their probability distribution.
-A higher entropy value indicates more complex local language patterns.
-The calculation formula is:
-$Entropy=-\sum_{i}^{N_{types}}p_{i}\log_{2}(p_{i})$ (8)
-where $p_{i}$ is the probability of occurrence for the $i^{th}$ 2-gram pattern.
-Since both TTR and Entropy are sensitive to text length, to ensure a relatively fair comparison of dataset complexity, we randomly sample 700 entries from the test set of each dataset for calculation(as the ATIS dataset only contains 700 instances).
-Appendix B Details for Reinforcement Learning
+Thus, we are more focused on enhancing such models’ capabilities and expand their broader applications.
+Deeper Study on LLM Capability Transfer
+Based on our experiments, the hybrid strategy effectively enhances super-tiny models’ generalization while almost entirely retaining LLMs’ capabilities in domain-specific tasks.
+In future research, we will conduct a deeper investigation and theoretical analysis of this potentially powerful methodology, building upon prior work to gain a better understanding of this hybrid strategy in transferring LLMs’ capabilities to super-tiny models.
+Cross-domain Optimal Compression Exploration
+It is evident that compressing models to 0.5B while retaining their capabilities across all domains is not feasible.
+The limitation boundary remains unclear.
+So, further research and exploration are required to better understand and define these limits.
+Appendix A Details for Reinforcement Learning
 Our reward design consists of two components, Format Reward($S_{format}$) and Answer Reward ($S_{answer}$):
@@
 Figure 6: An example of optimization of the original Answer Reward design by requiring explicit repetition of the answer within the CoT framework.
-Appendix C Prompt Compression
+Appendix B Prompt Compression
 To reduce inference costs, we remove unnecessary content in prompts, such as guidelines, format specifications, and few-shot examples used during large model invocation, while retaining essential elements: user queries, context, query-relevant knowledge, function lists, and brief function descriptions.
@@
 Figure 7: Original vs Compressed Prompt & Response.
-Appendix D Retriever Setup
+Appendix C Retriever Setup
 Initially, we select GTE-large [61], which is trained on the BERT-large [68], as our embedding model.
@@
 Subsequently, we switch to GTE-Qwen2-1.5B-instruct [61], a model trained on Qwen2-1.5B [71], which offers the advantage of fine-tuning for matching tasks through prompt modification.
-Next, utilizing this model with a customized prompt in Section D.1, we perform cosine similarity matching between the model-generated API descriptions and the JSON text of the entire API documentation, achieving a higher accuracy.
-During training, we optimize the API description model using rejection sampling, with implementation details provided in Section D.2.
+Next, utilizing this model with a customized prompt in Section C.1, we perform cosine similarity matching between the model-generated API descriptions and the JSON text of the entire API documentation, achieving a higher accuracy.
+During training, we optimize the API description model using rejection sampling, with implementation details provided in Section C.2.
 Finally, to achieve optimal inference acceleration, we train a classifier based on BERT-large [68] for top-1 recall using rejection sampling.
@@
 As illustrated in Table 4, this series of optimizations lead to a substantially improvement in API recall rates.
-Table 5: Comparison of retrieval performance of models: GTE-large model, GTE-Qwen2-1.5B-instruct model, optimized API description model + GTE-Qwen2-1.5B-instruct model.
+Table 4: Comparison of retrieval performance of different models: GTE-large model, GTE-Qwen2-1.5B-instruct model, optimized API description model + GTE-Qwen2-1.5B-instruct model.
 Extensibility implies that the API pool can be dynamically modified.
@@
 Extensibility ×
-D.1 Prompt for retrieval
+C.1 Prompt for retrieval
 {mdframed}
@@
 Query:
-D.2 Retriever Optimization
+C.2 Retriever Optimization
 Algorithm 1 Simplified API Description Validation Process
@@
 12:  end for
-Appendix E Training Details
-Table 6: Training Parameters in different phases.
+Appendix D Training Details
+Table 5: Training parameters of different knowledge transfer strategies.
 Supervised Fine-Tuning (SFT)
@@
 Logits Top-K $100$
-Table 7: Different KD results on Qwen2.5-0.5B model.
+Table 6: Ablation of KD on Qwen2.5-0.5B models.
 Model AR
@@
 Qwen2.5-0.5B (FKL) 94.65%
-[…]
-[…]
-[…]
+Qwen2.5-0.5B (AKL[$\mu=0.5$]) 94.48%
+Qwen2.5-0.5B (AKL[$\mu=0.7$]) 94.68%
+Qwen2.5-0.5B (AKL[$\mu=0.9$]) 94.85%
 The main training parameters are detailed in Table 6.
-During the training phase, we employ the AdamW optimizer [77], setting the model’s maximum sequence length to $5120$.
+During the training phase, we employ the AdamW optimizer [72] with the cosine learning rate schedule, setting the model’s maximum sequence length to $5120$.
 In the RL phase, we utilize the baseline variant of Reinforce++ [50] as the reinforcement learning algorithm, offering advantages in efficiency, conciseness, and training stability.
@@
 In the AKL phase, the threshold parameter $\mu$ is reduced from 1 to 0.5.
+When training with data containing CoT, we set the number of training epochs to 2 and use the average of the FKL loss and the CE loss.
 For pruning, we use the 0.5B-Instruct model as the base, implementing width/depth pruning strategies with a 20%/30% prune ratio, resulting in two models Qwen2.5-0.4B(0.35B)-Width and Qwen2.5-0.4B(0.35B)-Depth.
 These extremely compressed models are produced through a two-phase training process involving accuracy recovery training and knowledge distillation (KD).
-Comprehensive training details for the pruned models are available in Appendix I.
+Comprehensive training details for the pruned models are available in Appendix H.
 For quantization, we primarily experiment with three strategies on the 0.5B model trained with AKL: W8A16, W4A16, and W8A8.
@@
 In the evaluation phase, we employ SGLang [77] to generate data.
-Appendix F Computing resources
+Appendix E Computing resources
 We utilize 8 NVIDIA H20 GPUs for training and data generation processes.
@@
 Moreover, for the LLM prototype and LLM-as-a-Judge, generating data for every 10,000 entries takes approximately 1 hour.
-Appendix G Detailed Analysis for Experimental Results for KD
+Appendix F Detailed Analysis for Experimental Results for KD
 As demonstrated in Table 6, it is clear that balancing FKL and RKL is crucial for our task.
@@
 Overall, AKL demonstrates greater potential compared to RFT, achieving performance improvements with an appropriately chosen $\mu$ value.
-Appendix H Inference Details
+Appendix G Inference Details
 The inference framework utilizes TensorRT-LLM[67] v0.16.0 with enabled optimizations including paged attention, prefix cache, CUDA graph size set to 1024, and max token number set to 32768 for inference acceleration.
-Appendix I Pruning Details
-I.1 Offline One-shot Pruning
+Appendix H Pruning Details
+H.1 Offline One-shot Pruning
 Importance as Metric for Structured Pruning
@@
 The importance scores are first aggregated across batch and sequence dimensions through summation $\sum_{b,s}$, then normalized using L2 normalization within each subgroup to obtain global channel importance.
-Figure 8: Pruning Dependency in the Qwen2.5
+Figure 8: Width pruning dependency in the Qwen2.5 model.
 Optimal Width Pruning Configuration
@@
 This approach results in multiple width-pruned model configurations, as show in Table 7.
-We calculate channel importance scores $CI^{(i)}_{\text{embedding}}$ and $CI^{(j)}_{\text{inter}}$ using a calibration dataset consisting of 1024 randomly sampled sequences from WikiText-2.
+We calculate channel importance scores $CI^{i}_{\text{embedding}}$ and $CI^{j}_{\text{inter}}$ using a calibration dataset consisting of 1024 randomly sampled sequences from WikiText-2.
 For each configuration, we prune entire channel groups with the lowest CI scores and evaluate the resulting one-shot pruned models through WikiText-2 perplexity.
-Table 8 presents the optimal configurations for width pruning at rates of 20% and 30%, identified as Qwen2.5-0.4B-Width-2 and Qwen2.5-0.35B-Width-3, respectively.
-Table 8: Perplexity (PPL) on WikiText-2 for different model architectures with 20% and 30% width pruning.
+Table 7 presents the optimal configurations for width pruning at ratio of 20% and 30%, identified as Qwen2.5-0.4B-Width-2 and Qwen2.5-0.35B-Width-3, respectively.
+Table 7: Perplexity(PPL) evaluation on WikiText-2 comparing model architectures with one-shot width pruning at 20% and 30% pruning ratios.
 Model
@@
 Qwen2.5-0.35B-Width-5 640 4864 0.256 0.097 0.353 2395.52
-I.2 Recovery Training in Pruning
+H.2 Recovery Training in Pruning
 Recovery Training Dataset and Hyperparameters
@@
 We conduct training over 2 epochs, utilizing a batch size of 32, with each batch packed to accommodate sequences of up to 7168 tokens.
+Performance across Knowledge Assessment & Reasoning Tasks
+Table 8 summarizes the performance of pruned models across various downstream tasks, including MMLU[79] and CMMLU[80] for knowledge assessment, alongside ARC-Challenge[81] and HellaSwag[82] for commonsense reasoning evaluation, and GSM8k[83] for logical reasoning.
+Our experiments reveal distinct recovery patterns between width and depth pruning across general tasks.
+At a 20% pruning rate, width pruning achieves 90.5% of the performance relative to the unpruned baseline, while depth pruning retains only 78.4%.
+When increasing the pruning intensity to 30%, width pruning preserves 83.8% of the original performance.
+In contrast, depth pruning suffers catastrophic performance degradation, particularly on the GSM8K task, where accuracy drops from 45.11 to a merely 5.61.
 Table 8: Performance comparison of Qwen2.5 models after pruning.
 We report 5-shot performance on MMLU and CMMLU, 4-shot for GSM8k, zero-shot for ARC-Challenge, HellaSwag.
-* denotes that the baseline Qwen-2.5-0.5B utilizes the same instruct fine-tuning dataset for correction.
+* denotes that the baseline Qwen-2.5-0.5B utilizes the same instruct fine-tuning dataset for correction fine-tuning.
 Model
@@
 Qwen-2.5-0.35B-Depth 25.46 24.42 26.12 25.21 5.61 21.36
-Performance across Knowledge Assessment & Reasoning Tasks
-Table 9 summarizes the performance of pruned models across various downstream tasks, including MMLU and CMMLU for knowledge assessment, alongside ARC-Challenge, HellaSwag and Winogrande for commonsense reasoning evaluation.
-Our experiments reveal distinct recovery patterns between width and depth pruning across general tasks.
-At a 20% pruning rate, width pruning achieves 90.5% of the performance relative to the unpruned baseline, while depth pruning retains only 78.4%.
-When increasing the pruning intensity to 30%, width pruning preserves 83.8% of the original performance.
-In contrast, depth pruning suffers catastrophic performance degradation, particularly on the GSM8K task, where accuracy drops from 45.11 to a merely 5.61.
 Ablation of Pruning Ratio on Domain-Specific Task
@@
 These findings suggest that for small-scale models on domain-specific tasks, pruning rates exceeding 20% may significantly compromise accuracy.
-Table 10: Results with different pruning strategies on domain specific tasks.
+Table 9: Results in different pruning strategies on domain-specific tasks.
 Model AR
@@
 Qwen2.5-0.35B-Depth (KD) 93.75%
-Appendix J FLOPs Calculation
+Appendix I FLOPs Calculation
 For the calculation of FLOPs, we first define the notations symbolically:
@@
 We calculate FLOPs under $s_{\text{u}}=128$ tokens in the prefill phase and under $s_{\text{u}}=1$ token for each step in the decoding phase, with detailed results presented in Table 1.
-Appendix K Extensive Attempts on Generalization Enhancement via RL
-Table 11: Full results during the RL exploration
+Appendix J Extensive Attempts on Generalization Enhancement
+Table 10: In-domain and out-of-domain performance of Qwen2.5 models using different knowledge transfer strategies.
 Model AR
@@
 (Out-of-Domain)
-Crowdsourced Labelers 77.85% -
 BERT-based system 81.43% -
@@
 Qwen2.5-0.5B (RFT) 94.66% 90.62%
-Qwen2.5-0.5B (KD) 94.66% 90.62%
+Qwen2.5-0.5B (KD) 94.85% 90.53%
 Qwen2.5-7B (RL) 95.36% 94.80%
@@
 Qwen2.5-0.5B (RL) 90.76% 85.21%
+Knowledge Transfer using Qwen2.5-3B (RL)
+Qwen2.5-0.5B (RFT*) 93.82% 89.13%
+Qwen2.5-0.5B (RFT*+RL) 95.72% 90.09%
 Qwen2.5-0.5B (KD*) 88.49% 84.14%
 Qwen2.5-0.5B (KD*+RL) 95.55% 91.68%
-While attempting to apply reinforcement learning (RL), we conduct extensive experiments, with results shown in Appendix K.
-Initially, we find that directly training a 7B model with RL yields excellent results, surpassing the LLM Prototype and RFT-trained models.
-Subsequently, we attempt to train smaller models (3B and 0.5B) and discover that RL is ineffective on the 0.5B model, because the model tends to generate shorter or even meaningless Chain-of-Thought (CoT), leading to degraded output diversity as well as that without CoT.
-To address the issue above, we explore another approach, transferring out-of-domain capabilities into the 0.5B model via Knowledge Distillation (KD).
-However, we observe that the model struggles with format errors due to underfitting.
-Ultimately, after applying RL training on this model, we notice an impressive improvement on model performance, with both an absolute 1% AR increase on in-domain/out-of-domain datasets compared to the 0.5B RFT-trained model.
+Knowledge Transfer using Qwen2.5-7B (RL)
+Qwen2.5-0.5B (RFT*) 95.29% 90.28%
+Qwen2.5-0.5B (RFT*+RL) 95.94% 92.21%
+Qwen2.5-0.5B (KD*) 95.32% 90.59%
+Qwen2.5-0.5B (KD*+RL) 95.99% 92.34%
+Figure 9: Response lengths of different models during RL training
+Table 11: Response lengths of different models, represented as mean $\pm$ standard deviation.
+Model Response Length
+QwQ-32B (ICL) 316.8 $\pm$ 111.4
+Qwen2.5-7B (RL) 86.4 $\pm$ 12.9
+Qwen2.5-3B (RL) 61.7 $\pm$ 9.4
+Qwen2.5-0.5B (KD*) 86.3 $\pm$ 12.6
+Qwen2.5-0.5B (KD*+RL) 85.3 $\pm$ 10.8
+Upon discovering that using RFT or KD alone leads to significant degradation in out-of-domain performance for super-tiny models, we conduct extensive experiments to enhance generalization capabilities, as illustrated in Appendix J.
+Initially, we prioritize Reinforcement Learning (RL) for this task.
+As expected, 7B RL-trained models achieve excellent results, surpassing the LLM Prototype and RFT-trained models.
+Subsequently, we attempt direct RL training on smaller models, such as 3B and 0.5B.
+The results are less favorable this time.
+While 3B RL-trained models exhibit considerable generalization capabilities, 0.5B RL-trained models show no improvement in generalization, underperforming even compared to 0.5B RFT-trained models.
+Through observations and analyses, we conclude that this is because models of such small size tend to generate low-quality Chain-of-Thought (CoT), as shown in Figure 9, leading to degraded output diversity, akin to scenarios where CoT is not used.
+To address the challenge of 0.5B models’ inabilities to generate high-quality CoT, we propose a hybrid strategy.
+Given the favorable results from 7B RL-trained models, we decide to use 7B RL-trained models as our starting point.
+Then, we transfer out-of-domain capabilities and the abilities of generating high-quality CoT from 7B RL-trained models to 0.5B models via Rejection sampling Fine-Tuning (RFT) or knowledge distillation (KD), which are denoted as RFT* and KD*, respectively.
+Experimental results indicate that this knowledge transfer indeed outperforms direct RL training on 0.5B models in enhancing generalization and preserving the abilities of generating high-quality CoT.
+So far, 0.5B KD* models posses the abilities to generate high-quality CoT, but in terms of out-of-domain Achievable Rate (AR), there is no considerable improvement compared to 0.5B RFT-trained models.
+Consequently, we conduct another round of RL training on 0.5B KD* models with data generated by the LLM prototype for enhancing generalization, similar to how we initially improve 7B models’ generalization with RL.
+Ultimately, experimental results present that after the second RL training, the 0.5B super-tiny models exhibit impressive improvement on model performance, with both an absolute increase of over 1% AR increase on in-domain and out-of-domain datasets compared to 0.5B RFT-trained models.
+Table 12: Ablation of knowledge transfer strategies on Qwen2.5-0.5B models.
+Model AR
+AR
+(Out-of-Domain)
+Qwen2.5-0.5B (RFT*) 95.29% 90.28%
+Qwen2.5-0.5B (KD*[FKL]) 95.08% 90.19%
+Qwen2.5-0.5B (KD*[0.5FKL+0.5CE]) 95.32% 90.59%
+Qwen2.5-0.5B (KD*[AKL]) 93.93% 88.37%
+Qwen2.5-0.5B (KD*[0.5AKL+0.5CE]) 95.82% 90.85%
+Qwen2.5-0.5B (RFT*+RL) 95.94% 92.21%
+Qwen2.5-0.5B (KD*[FKL]+RL) 95.66% 92.21%
+Qwen2.5-0.5B (KD*[0.5FKL+0.5CE]+RL) 95.99% 92.34%
+Qwen2.5-0.5B (KD*[AKL]+RL) 95.85% 92.10%
+Qwen2.5-0.5B (KD*[0.5AKL+0.5CE]+RL) 96.10% 92.03%
+Ablation of Different Teacher Models
+As illustrated in Appendix J, a teacher model with inherently superior capabilities is crucial for enhancing student model performance, particularly regarding out-of-domain capabilities.
+A better teacher model not only leads to better performance but also ensures greater stability in the student models’ results, driven by the generation of higher-quality CoT.
+Response Lengths of Different Reasoning Models
+As can be seen in Table 11, the average response lengths of our models are more than three times shorter than that of QwQ-32B.
+For direct RL training, longer response lengths appear to yield better results.
+However, this conclusion is not entirely valid.
+For instance, in the case of KD* models, there is almost no change in response lengths before and after RL, yet there is a significant difference in out-of-domain performance.
+Ablation of Knowledge Transfer Strategies
+During the knowledge transfer stage from 7B RL-trained Qwen2.5 models, we investigate the effectiveness of different knowledge transfer strategies, as illustrated in Appendix J.
+Apart from the KD*[AKL] strategy, which encounters a certain degree of underfitting, other experimental results demonstrate the robustness of different knowledge transfer strategies.
+Specifically, utilizing a loss with Cross Entropy (CE) is preferable to not incorporating CE during KD, as the high-quality CoT generated by 7B RL-trained models facilitates faster convergence with CE loss.
+Conversely, when the teacher model generates low-quality CoT, using CE alone, equivalent to RFT, can lead to overfitting, as evidenced in Appendix J, where the 3B RL-trained model as the teacher model exhibits poorer performance.
+Appendix K Experiments on Public Datasets
+Table 13: Performance of different models on BFCL and xlam-function-calling-60k datasets.
+BFCL In-domain
+Model Overall Non-live Live Acc
+Qwen2.5-0.5B 55.14 62.29 47.98 55.88
+Hammer2.0-0.5B [85] 59.31 67.00 51.62 /
+Qwen2.5-72B 83.55 88.13 78.98 79.47
+Qwen2.5-7B (RL) 81.85 86.27 77.42 78.55
+Qwen2.5-0.5B (RL) 55.15 64.48 45.82 65.68
+Qwen2.5-0.5B (RFT) 56.64 64.50 48.78 74.08
+Qwen2.5-0.4B-Depth (RFT) 42.27 54.56 29.98 71.55
+Qwen2.5-0.4B-Width (RFT) 42.60 56.17 29.02 72.98
+Qwen2.5-0.5B-W8A16 (RFT) 56.42 64.42 48.41 74.38
+Qwen2.5-0.5B-W4A16 (RFT) 49.54 58.52 40.56 72.23
+Qwen2.5-0.5B-W8A8 (RFT) 55.68 62.88 48.48 74.06
+Knowledge Transfer using Qwen2.5-7B (RL)
+Qwen2.5-0.5B (KD*[FKL]) 59.44 69.06 49.81 72.97
+Qwen2.5-0.5B (KD*[0.5FKL+0.5CE]) 54.28 62.00 46.56 66.55
+Qwen2.5-0.5B (KD*[AKL]) 58.45 67.15 49.74 71.90
+Qwen2.5-0.5B (KD*[0.5AKL+0.5CE]) 58.97 68.12 49.81 71.70
+Qwen2.5-0.5B (KD*[FKL]+RL) 60.42 71.02 49.81 75.87
+Qwen2.5-0.5B (KD*[0.5FKL+0.5CE]+RL) 59.84 70.00 49.67 77.02
+Qwen2.5-0.5B (KD*[AKL]+RL) 60.50 69.56 51.44 76.38
+Qwen2.5-0.5B (KD*[0.5AKL+0.5CE]+RL) 60.79 69.46 52.11 76.68
+To validate the effectiveness of our proposed method, we conduct experiments on publicly available datasets.
+We randomly partition 90% of the xlam-function-calling-60k dataset [69] for training, with the remaining 10% serving as an in-domain test set.
+For the training set, instead of utilizing the provided labels, we employ Qwen2.5-72B-Instruct [11] to generate reference labels.
+Furthermore, we design a judge based on QwQ-32B [44] to assess the accuracy of the labels generated by the LLM.
+This process can be viewed as prototyping.
+Given that function calling is a standardized task, we use the official function call prompt template instead of developing a specialized agent for answer generation.
+The design of our judge is inspired by the Format Checker and Semantic Checker proposed in [69].
+Through this procedure, we ultimately curate 50,201 valid query-tools-answers pairs from the initial 54k training samples for subsequent training.
+Due to the complex function call formats within this dataset, we redesign the reward mechanism for Reinforcement Learning (RL). The reward is defined as:
+$Reward=\alpha\cdot Reward_{format}+\beta\cdot Reward_{answer}$ (11)
+where $Reward_{format}$ is 1 if and only if all the following conditions are met, and 0 otherwise: (1) The model’s output contains exactly one "<think>…</think>" block; (2) The function call within the answer can be parsed by the official Qwen prompt template and can be loaded as a JSON object; (3) The function name in the answer exists within the tools provided in the prompt, and the parameter types conform to their definitions, which is implemented using Pydantic[86].
+For $Reward_{answer}$, we first check whether the number of functions in the model’s output is equivalent to that of the ground truth: If not, $Reward_{answer}$ is 0; Otherwise, $Reward_{answer}$ is defined as follows:
+$Reward_{i}=\begin{cases}\frac{\lvert P_{i,pred}\cap P_{i,label}\rvert}{\lvert P_{i,pred}\cup P_{i,label}\rvert},&\text{if }\lvert P_{i,pred}\cup P_{i,label}\rvert>0\\ 1,&\text{otherwise}\end{cases}$ (12)
+$Reward_{answer}=\frac{1}{N}\sum Reward_{i}$ (13)
+where $N$ is the number of function calls, and $P_{i}$ denotes the set of parameters for the $i$-th function call.
+During training, we set $\alpha=1$ and $\beta=2$.
+The baseline variant of Reinforce++ [50] is selected as our reinforcement learning method.
+We evaluate the trained model on both in-domain and out-of-domain dataset.
+For in-domain evaluation, we utilize the aforementioned test set with official answers as the ground truth.
+A test instance is considered correct only if the model’s output exactly matches the answer (i.e., identical function names and key-value pairs for arguments); all other cases are deemed incorrect.
+Accuracy is calculated based on this criterion.
+Additionally, we employ the Berkeley Function Call Leaderboard (BFCL)[70] for out-of-domain evaluation.
+The final experimental results are presented in Table 13.
+The experimental results indicate that the 0.5B trained model achieves in-domain accuracy comparable to the Qwen2.5-72B model.
+Furthermore, it demonstrates a substantial improvement over Qwen2.5-0.5B-Instruct in out-of-domain scenarios.
+We also compare our method with Hammer2.0-0.5B[85], the previous SOTA 0.5B model, which is also fine-tuned from Qwen2.5-0.5B-Instruct8.
+Our approach achieves superior performance, demonstrating the effectiveness of our proposed method.
+In terms of quantization, experimental results demonstrate that applying W8A8 and W8A16 strategies results in negligible loss compared to the original model, while using the W4A16 strategy leads to significant loss in out-of-domain scenarios.
+Regarding pruning, all pruned models exhibit substantial losses in out-of-domain scenarios, yet own relatively minor loss in in-domain scenarios.